### v1

# 7491, Trondheim, Norway

Abstract
In this paper, we propose criteria for unboundedness of the images of set-valued mappings having closed graphs in Euclidean spaces. We focus on mappings whose domains are non-closed or whose values are connected. These criteria allow us to see structural properties of solutions in vector optimization, where solution sets can be considered as the images of solution mappings associated to specific scalarization methods. In particular, we prove that if the domain of a certain solution mapping is non-closed, then the weak Pareto solution set is unbounded. Furthermore, for a quasi-convex problem, we demonstrate two criteria to ensure that if the weak Pareto solution set is disconnected then each connected component is unbounded.
Keywords:
Set-valued mapping closed graph unboundedness disconnectedness vector optimization scalarization quasi-convex function
MSC:
90C33 90C3154C60
1 Introduction
Studying structure of solutions plays an important role in vector optimization theory, see, for example (lucbook, , Chapter 6) and (ehrgott2005multicriteria, , Chapter 3). The most effective method for such studying is considering solution sets as the images of solution mappings based on scalarization techniques.
It is well-known that the image of a connected set under a continuous single-valued mapping is connected. Naccache in naccache1978connectedness , Warburton in warburton1983quasiconcave , and Hiriart-Urruty in hiriart1985images extended the result for a set-valued mapping having connected values, provided that the mapping is upper (or lower) semi-continuous, see Theorem 2.1. This criterion opened a door to study connectedness of solution sets of not only vector optimization problems but also of set optimization problems, vector variational inequalities, vector equilibrium problems. There are numerous works related to this topic, we can only list a few of them, for example luc1987connectedness ; sun1996connectedness ; gong2007connectedness ; yen2011monotone ; han2018existence ; sharma2022connectedness ; peng2019connectedness ; han2022connectedness ; anh2022connectedness . Almost all of the obtained results require generalized convexity or monotonicity assumptions to ensure connected-valuedness, and require boundedness of the constraint sets or certain coercivity conditions to guarantee semi-continuity of solution mappings.
The aim of this paper is to look for tools that can be applied to investigate unboundedness and disconnectedness of solutions to problems governed by unbounded data in optimization theory.
In particular, the first part is dedicated to studying unboundedness of the images of set-valued mappings having closed graphs. In Theorem 3.1, we prove that if a mapping has non-closed domain then its image is unbounded. We next study the relationship between unboundedness and disconnectedness of the image of a mapping having connected values. In Theorem 4.1, we introduce two criteria to ensure unboundedness of all connected components of the image.
In the second part, we apply the above results to vector optimization based on two solution mappings associated to two scalarization methods. We obtain new properties related to unboundedness of the (weak) Pareto solution set of a problem in both cases: with a generalized convexity assumption and without imposing generalized convexity. For the second case, in Theorem 5.2 we prove that if the domain of a solution mapping between two aforementioned ones is non-closed then the weak Pareto solution set is unbounded.
For a problem governed by semi-strictly quasi-convex functions, to ensure disconnectedness of all components of the weak Pareto solution set, Hoa, Huy, Phuong and Yen in (hoa2007unbounded, , Theorem 3.2) used a condition which says that the number of components of the set is finite. In Theorem 6.1, we introduce two other criteria for quasi-convex problems. As a result, when the objective functions are strictly quasi-convex, the Pareto solution set and weak Pareto solution set are coincident, and if this set is disconnected then all the components are unbounded.
The remaining part of this paper consists of five sections. Section 2 gives some definitions, notations, and auxiliary results on set-valued mappings and connected sets. Sections 3 and 4 establish results concerning unboundedness of the images of set-valued mappings having non-closed domains and having connected values, respectively. Results on unboundedness of solutions to vector optimization problems are shown in the last two sections, namely Sections 5 and 6.
2 Preliminaries
Let $X$ be a nonempty subset of $\operatorname{\mathbb{R}}^{m}$. We consider set-valued mapping $S$ from $X$ to $\operatorname{\mathbb{R}}^{n}$, $S:X\rightrightarrows\operatorname{\mathbb{R}}^{n}$. The domain of $S$ is the set of all elements $u$ in $X$ such that $S(u)$ is non-empty:
$\operatorname{dom}S\coloneqq\{u\in X:S(u)\neq\emptyset\}$.
The union of the images (or values) $S(u)$, when $u$ ranges over $U\subseteq X$, is the image of $U$ under the mapping $S$:
$S(U)\coloneqq\cup_{u\in U}S(u).$
The image of $S$ is the set $S(X)=S(\operatorname{dom}S)$. If its graph
$\operatorname{gph}S\coloneqq\{(u,v)\in X\times\operatorname{\mathbb{R}}^{n}:v\in S(u)\}$
is a closed subset of $X\times\operatorname{\mathbb{R}}^{n}$ then one says that $S$ has closed graph. If $S$ has closed graph then $S(u)$ is a closed set for all $u\in X$. The inverse mapping of $S$, denoted by $S^{-1}:\operatorname{\mathbb{R}}^{n}\rightrightarrows X$, is defined by
$S^{-1}(v)\coloneqq\{u\in X:v\in S(u)\},\ v\in\operatorname{\mathbb{R}}^{n}.$
If $\operatorname{\mathcal{A}}$ is a subset in $\operatorname{\mathbb{R}}^{n}$, then the inverse image of $\operatorname{\mathcal{A}}$ by $S$ is the following set
$S^{-1}(\operatorname{\mathcal{A}})=\{u\in X:S(u)\cap\operatorname{\mathcal{A}}\neq\emptyset\}.$
The mapping $S$ is locally bounded at $\bar{u}$ if there exists an open neighborhood $U$ of $\bar{u}$ such that $S(U\cap X)$ is bounded. If $S$ is locally bounded at every $u\in X$ then $S$ is said to be locally bounded on $X$. The mapping $S$ is upper semi-continuous at $\bar{u}\in X$ if for any open set $V\subseteq\operatorname{\mathbb{R}}^{n}$ such that $S(\bar{u})\subseteq V$ there exists an open neighborhood $U$ of $\bar{u}$ such that $S(u)\subseteq V$ for all $u\in U$. If $S$ is upper semi-continuous at every $u\in X$ then $S$ is said to be upper semi-continuous on $X$.
From (AuslenderTeboulle2003AsymptoticConesFunctionsOptimizationVariationalInequalitiesbook, , Theorem 1.4.1), one has the following lemma which will be used in the proof of Theorem 4.1.
Lemma 1
Assume that $S$ has closed graph. If $S$ is locally bounded at $\bar{u}$ for some $\bar{u}$ in $\operatorname{dom}S$, then $S$ is upper semi-continuous at $\bar{u}$.
Let $Y$ be a nonempty subset of $\operatorname{\mathbb{R}}^{n}$. One says that $Y$ is connected if there do not exist two nonempty disjoint subsets $A,B$ of $Y$ with $Y\subseteq A\cup B$ and two open subsets $U,V$ in $\operatorname{\mathbb{R}}^{n}$ such that $A\subseteq U$, $B\subseteq V$ and $U\cap V=\emptyset$. Otherwise $Y$ is called disconnected. A nonempty subset $A\subseteq Y$ is said to be a connected component of $Y$ if $A$ is connected and it is not a proper subset of any connected subset of $Y$. Recall that when $Y$ is connected, if the set $A\subseteq Y$ is closed and open in $Y$ then $A=Y$. Moreover, if $Y$ is connected then the union $Y\cup\{y\}$, where $y$ belongs to the closure $\operatorname{cl}(Y)$, is also connected.
One says that the set-valued mapping $S$ has connected (convex, bounded) values if $S(x)$ is connected (convex, bounded) for every $x$ on the domain of $S$.
In order to be convenient for readers to compare our developments with the before-mentioned result by Naccache, Warburton, and Hiriart-Urruty, we recall the following theorem; its proofs can be found in (naccache1978connectedness, , Lemma 2.1), (warburton1983quasiconcave, , Theorem 3.1), or (hiriart1985images, , Theorem 3.1):
Theorem 2.1
Assume that $X$ is connected and $S$ has connected values. If $S$ is upper semi-continuous on $X$, then the image $S(X)$ is connected.
3 Set-valued mappings having non-closed domains
Let $A$ be a nonempty subset of $\operatorname{\mathbb{R}}^{m}$. We want to emphasize that, throughout the present section, the set-valued mapping
$\Phi:A\rightrightarrows\operatorname{\mathbb{R}}^{n},\ a\mapsto\Phi(a),$
has closed graph. We denote by $\operatorname{ri}A$ its relative interior.
The proofs of the main results in the present and next sections are based on the following lemma:
Lemma 2
If $\operatorname{\mathcal{A}}$ is a bounded connected component of $\Phi(A)$, then $\Phi^{-1}(\operatorname{\mathcal{A}})$ is closed. Furthermore, if $\Phi(A)$ is a bounded, then $\operatorname{dom}\Phi$ is closed.
Proof
Let $\operatorname{\mathcal{A}}$ be a bounded connected component of $\Phi(A)$. To prove that $\Phi^{-1}(\operatorname{\mathcal{A}})$ is closed, we suppose that the sequence $\{a^{k}\}\subseteq\Phi^{-1}(\operatorname{\mathcal{A}})$ converges to $\bar{a}\in\operatorname{\mathbb{R}}^{m}$ and need to prove $\bar{a}\in\Phi^{-1}(\operatorname{\mathcal{A}})$. Indeed, let $\{x^{k}\}\subseteq\operatorname{\mathcal{A}}$ be a sequence such that $x^{k}\in\Phi(a^{k})$ for every $k\in\operatorname{\mathbb{N}}$. By the boundedness of $\operatorname{\mathcal{A}}$, we can assume that $x^{k}\to\bar{x}$. Since the graph of $\Phi$ is closed, $(\bar{a},\bar{x})$ belongs to $\operatorname{gph}\Phi$, thus $\bar{x}\in\Phi(\bar{a})$. As $\operatorname{\mathcal{A}}$ is connected, the union $\operatorname{\mathcal{A}}\cup\{\bar{x}\}$ is connected. This means that $\bar{x}\in\operatorname{\mathcal{A}}$. We thus get $\bar{x}\in\operatorname{\mathcal{A}}\cap\Phi(\bar{a})\neq\emptyset$, and then $\bar{a}\in\Phi^{-1}(\operatorname{\mathcal{A}})$ by definition. Therefore, the inverse image $\Phi^{-1}(\operatorname{\mathcal{A}})$ is closed.
The second conclusion is proven similarly. ∎
Theorem 3.1
Assume that $\operatorname{dom}\Phi$ is non-closed. Then, $\Phi(A)$ is unbounded. Moreover, if additionally $A$ is convex and $\Phi(\operatorname{ri}A)$ is nonempty then the latter set is also unbounded.
Proof
By the non-closedness of $\operatorname{dom}\Phi$ and the second conclusion in Lemma 2, we conclude straightforwardly that $\Phi(A)$ is unbounded. We only need to prove the second claim.
Now let $A$ be convex. Suppose that $\Phi(\operatorname{ri}A)$ is nonempty but on the contrary the set is bounded. Let $\overline{\Sigma}$ be the closure of the following set:
$\Sigma\coloneqq\left(\operatorname{gph}\Phi\right)\cap\left(\operatorname{ri}A\times\operatorname{\mathbb{R}}^{n}\right)=\{(a,x)\in\operatorname{ri}A\times\operatorname{\mathbb{R}}^{n}:x\in\Phi(a)\}$.
Clearly, one has $\Sigma\subseteq\operatorname{gph}\Phi$. Since $\operatorname{gph}\Phi$ is closed, the closure of $\Sigma$ is also a subset of $\operatorname{gph}\Phi$, i.e. $\overline{\Sigma}\subseteq\operatorname{gph}\Phi$. Let $\pi:\operatorname{\mathbb{R}}^{m+n}\to\operatorname{\mathbb{R}}^{n}$ be the projection defined by $\pi(t_{1},\dots,t_{m},t_{m+1},\dots,t_{m+n})=(t_{m+1},\dots,t_{m+n})$. Clearly, the image of $\overline{\Sigma}$ under $\pi$ is bounded and
$\Phi(\operatorname{ri}A)\subseteq\pi(\overline{\Sigma})\subseteq\Phi(A).$
We consider a new mapping $\Phi^{\prime}$ defined by
$\Phi^{\prime}:A\rightrightarrows\operatorname{\mathbb{R}}^{n},\ \Phi^{\prime}(a)\coloneqq\pi(\overline{\Sigma})\cap\Phi(a).$
One can check that the graph of $\Phi^{\prime}$ coincides with $\overline{\Sigma}$. Hence, $\Phi^{\prime}$ also has closed graph. By boundedness of $\pi(\overline{\Sigma})$, $\Phi^{\prime}(A)=\pi(\overline{\Sigma})$ is also bounded. Applying the second conclusion in Lemma 2 to $\Phi^{\prime}$, $\operatorname{dom}\Phi^{\prime}$ is closed. Furthermore, it is not difficult to see the following inclusions:
$(\operatorname{ri}A)\cap\operatorname{dom}\Phi\ \subseteq\ \operatorname{dom}\Phi^{\prime}\ \subseteq\ \operatorname{dom}\Phi.$
By convexity of $A$ and density of $(\operatorname{ri}A)\cap\operatorname{dom}\Phi$ in $\operatorname{dom}\Phi$, $\operatorname{dom}\Phi^{\prime}$ is also dense in $\operatorname{dom}\Phi$. It follows that $\operatorname{dom}\Phi$ is closed. This is a contradiction. ∎
Remark 1
Let $\operatorname{\mathcal{A}}$ be a connected component of $\Phi(A)$. If $\Phi^{-1}(\operatorname{\mathcal{A}})$ is non-closed, thanks to Lemma 2, $\operatorname{\mathcal{A}}$ is unbounded.
[…]
Figure 1: The image of $\Theta_{1}$ has two unbounded components.
Example 1
From Example 6 we consider the set-valued mapping that is the solution mapping $\Theta_{1}:\Delta\rightrightarrows\operatorname{\mathbb{R}}^{2}$ of the there given vector optimization problem:
$\Theta_{1}(\xi_{1})=\left\{\begin{array}[]{cl}\left\{\Big{(}\pm\sqrt{\frac{1-\xi_{1}}{\xi_{1}}},\frac{\xi_{1}}{1-\xi_{1}}\Big{)}\right\}&\ \ \hbox{ if }\ \xi_{1}\in(0,1),\\ \emptyset&\ \ \hbox{ if }\ \xi_{1}\in\{0,1\}.\\ \end{array}\right.$ (3.1)
We can see that the graph of $\Theta_{1}$ is closed and that the domain of $\Theta_{1}$ is the open interval $(0,1)$ which is non-closed. According to Theorem 3.1, the image of $\Theta_{1}$ is unbounded. In particular, we have
$\Theta_{1}([0,1])=\big{\{}(x_{1},x_{2})\in\operatorname{\mathbb{R}}^{2}:x_{1}^{2}x_{2}=1\big{\}}=\operatorname{\mathcal{A}}\dot{\cup}\operatorname{\mathcal{B}},$ (3.2)
where
$\operatorname{\mathcal{A}}\coloneqq\Big{\{}(x_{1},x_{2})\in\operatorname{\mathbb{R}}^{2}:x_{2}>0,x_{1}\sqrt{x_{2}}=-1\Big{\}}$
and
$\operatorname{\mathcal{B}}\coloneqq\Big{\{}(x_{1},x_{2})\in\operatorname{\mathbb{R}}^{2}:x_{2}>0,x_{1}\sqrt{x_{2}}=1\Big{\}}.$
Clearly, $\Theta_{1}^{-1}(\operatorname{\mathcal{A}})=\Theta_{1}^{-1}(\operatorname{\mathcal{B}})=\operatorname{ri}\Delta$ is non-closed. Hence, $\Theta_{1}([0,1])$ has two unbounded connected components $\operatorname{\mathcal{A}}$ and $\operatorname{\mathcal{B}}$, illustrated in Figure 1, whose inverse images are non-closed.
4 Set-valued mappings having connected values
It is important to underline that, throughout the section, we consider another set-valued mapping
$\Psi:A\rightrightarrows\operatorname{\mathbb{R}}^{n},\ a\mapsto\Psi(a),$
where $A\subseteq\operatorname{\mathbb{R}}^{m}$ is nonempty, and $\Psi$ has closed graph and connected values.
In Theorem 2.1, under upper semi-continuity condition, the image of a connected set through $\Psi$ is connected. Consequently, if the image is disconnected then either $\Psi$ is not upper semi-continuous or the domain of $\Psi$ is disconnected. In this section, we focus on the relationship between disconnectedness of $\Psi(A)$ and its unboundedness.
Lemma 3
If $\operatorname{\mathcal{A}}$ and $\operatorname{\mathcal{B}}$ are different connected components of $\Psi(A)$, then
$\Psi^{-1}(\operatorname{\mathcal{A}})\cap\Psi^{-1}(\operatorname{\mathcal{B}})=\emptyset.$ (4.1)
Proof
Suppose that $\operatorname{\mathcal{A}}$ and $\operatorname{\mathcal{B}}$ are different connected components of $\Psi(A)$. Assume on the contrary that there exists $a\in\Psi^{-1}(\operatorname{\mathcal{A}})\cap\Psi^{-1}(\operatorname{\mathcal{B}}).$ It follows that $\Psi(a)\cap\operatorname{\mathcal{A}}\neq\emptyset$ and $\Psi(a)\cap\operatorname{\mathcal{B}}\neq\emptyset$. Note that $\Psi(a)$ is connected. Since $\operatorname{\mathcal{A}},\operatorname{\mathcal{B}}$ are connected components of $\Psi(A)$, one has $\Psi(a)\subseteq\operatorname{\mathcal{A}}$ and $\Psi(a)\subseteq\operatorname{\mathcal{B}}$. This yields
$\Psi(a)\subseteq\operatorname{\mathcal{A}}\cap\operatorname{\mathcal{B}}\neq\emptyset,$
which contradicts the assumption that $\operatorname{\mathcal{A}}$ and $\operatorname{\mathcal{B}}$ are two different connected components. Thus (4.1) is proved. ∎
The next two theorems provide criteria for unboundedness of all connected components of the image of a mapping having connected values. The technique used to prove these theorems is generalized from yen2011monotone ; hieu2021disconnectedness . The idea of these theorems is the following observation: we are interested in unboundedness, hence we ignore the unbounded values of $\Psi$ and focus on the points in $\operatorname{dom}\Psi$ having bounded values. Here, we denote by $\mathcal{O}$ the set
$\mathcal{O}=\{a\in\operatorname{dom}\Psi:\Psi(a)\text{ is bounded}\}$.
Theorem 4.1
Assume that $\operatorname{dom}\Psi$ is connected and that $\Psi$ is locally bounded or upper semi-continuous on $\mathcal{O}$. Then, if $\Psi(A)$ is disconnected, all its connected components are unbounded.
Proof
Recall that the mapping $\Psi$ has closed graph. Thanks to Lemma 1, local boundedness of $\Psi$ implies its upper semi-continuity on $\mathcal{O}$. Hence, we only need to consider the case that $\Psi$ is upper semi-continuous on $\mathcal{O}$.
Suppose that $\Psi(A)$ is disconnected and $\operatorname{\mathcal{A}}$ is a connected component of $\Psi(A)$. On the contrary, we assume that $\operatorname{\mathcal{A}}$ is bounded. Let $\operatorname{\mathcal{B}}$ is another connected component of $\Psi(A)$. Then there exists an open set $V$ containing $\operatorname{\mathcal{A}}$ such that $V\cap\operatorname{\mathcal{B}}=\emptyset$. Recall from Lemma 3 that the two sets $\Psi^{-1}(\operatorname{\mathcal{A}})$ and $\Psi^{-1}(\operatorname{\mathcal{B}})$ are disjoint.
By the closedness of the graph of $\Psi$ and Lemma 2, $\Psi^{-1}(\operatorname{\mathcal{A}})$ is closed. We now point out that $\Psi^{-1}(\operatorname{\mathcal{A}})$ is open in $\operatorname{dom}\Psi$. Indeed, let $a$ be an arbitrary point in $\Psi^{-1}(\operatorname{\mathcal{A}})$, then $\emptyset\neq\Psi(a)\subseteq\operatorname{\mathcal{A}}$ is bounded. It follows that $a\in\mathcal{O}$. By the assumption, $\Psi$ is upper semi-continuous at $a$. With the open set $V$ defined above, there is an open set $U$ in $\operatorname{dom}\Psi$ such that $\Psi(a^{\prime})\subseteq V$ for all $a^{\prime}\in U$. Hence, the openness of $\Psi^{-1}(\operatorname{\mathcal{A}})$ is proved.
From above, $\Psi^{-1}(\operatorname{\mathcal{A}})$ is both open and closed in the connected set $\operatorname{dom}\Psi$. It follows that $\Psi^{-1}(\operatorname{\mathcal{A}})=\operatorname{dom}\Psi$, while $\operatorname{dom}\Psi$ contains $\Psi^{-1}(\operatorname{\mathcal{B}})\neq\emptyset$. This is impossible because of $\Psi^{-1}(\operatorname{\mathcal{A}})\cap\Psi^{-1}(\operatorname{\mathcal{B}})=\emptyset$. Thus, $\operatorname{\mathcal{A}}$ must be unbounded. ∎
To illustrate above theorem, we consider an example coming from linear fractional vector optimization.
Example 2
Consider the solution mapping of a linear fractional bi-criteria optimization problem in (hoa2005parametric, , Section 4), which has the feasible set
$K=\{(x_{1},x_{2})\in\operatorname{\mathbb{R}}^{2}:x_{1}\geq 0,x_{2}\geq 0,x_{1}+x_{2}\geq 1\}$
and the objective functions
$f_{1}(x)=\frac{x_{1}+1}{2x_{1}+x_{2}},\ \ f_{2}(x)=\frac{-x_{1}-2}{x_{1}+x_{2}}.$
$x_{1}$$x_{2}$$\xi_{1}=\frac{1}{2}$$\frac{2}{3}>\xi_{1}>\frac{1}{2}$$\xi_{1}=\frac{2}{3}$$\xi_{1}\in[0,\frac{1}{2})$
Figure 2: The image of $\Theta_{1}$ has two unbounded components.
The solution mapping $\Theta_{1}$ of the problem (see Section 5.1) can be written as follows:
$\Theta_{1}(\xi_{1})=\left\{\begin{array}[]{cl}\{(1,0)\}&\ \ \hbox{ if }\ 0\leq\xi_{1}<\frac{1}{2},\\ [1,+\infty)\times\{0\}&\ \ \hbox{ if }\ \ \ \ \xi_{1}=\frac{1}{2},\\ \left\{\Big{(}\frac{2-3\xi_{1}}{2\xi_{1}-1},2\Big{)}\right\}&\ \ \hbox{ if }\ \frac{1}{2}<\xi_{1}<\frac{2}{3},\\ \{0\}\times[2,+\infty)&\ \ \hbox{ if }\ \ \ \ \xi_{1}=\frac{2}{3},\\ \emptyset&\ \ \hbox{ if }\ \frac{2}{3}<\xi_{1}\leq 1.\\ \end{array}\right.$
The domain of $\Theta_{1}$ is the interval $[0,\frac{2}{3}]$ which is connected. Moreover, the set $\mathcal{O}$ is $[0,\frac{1}{2})\cup(\frac{1}{2},\frac{2}{3})$ and $\Theta_{1}$ is single-valued and continuous on $\mathcal{O}$. Hence, $\Theta_{1}$ satisfies the condition in above theorem. The image has two connected components, both of them are unbounded.
The below theorem proposes another criterion to replace connectedness of $\operatorname{dom}\Psi$ in Theorem 4.1.
Theorem 4.2
Assume that $A$ is connected, $\mathcal{O}$ is open in $A$, and $\Psi$ is upper semi-continuous or locally bounded on $\mathcal{O}$. Then, if $\Psi(A)$ is disconnected, all its connected components are unbounded.
Proof
The proof is similar to the proof of Theorem 4.1. Note that the openness of $\Psi^{-1}(\operatorname{\mathcal{A}})$ in $\operatorname{dom}\Phi$ is proved under the assumptions that $\mathcal{O}$ is open in $A$ and $A$ is connected.∎
The assumption related to openness of $\mathcal{O}$ and local boundedness or upper semi-continuity of $\Psi$ on $\mathcal{O}$ in Theorem 4.2 seems heavily in general. However, it is satisfied for several cases such as solution mappings in strictly quasi-convex vector optimization (see Lemma 6), linear fractional vector optimization, or monotone affine/general vector inequality yen2011monotone ; hieu2021disconnectedness .
We consider the following example that comes from monotone affine vector inequalities in which the assumption in Theorem 4.1 related to connectedness of the domain is not satisfied but the assumptions in Theorem 4.2 are fulfilled.
Example 3
We next consider the bi-criteria case of the monotone affine vector inequality given in the proof of (hieu2019numbers, , Theorem 3.2) whose feasible set is $K=\operatorname{\mathbb{R}}^{2}$, and the two mappings are
$F_{1}(x)=\begin{bmatrix}x_{2}+1\\ -x_{1}+1\par\end{bmatrix},\ F_{2}(x)=\begin{bmatrix}-x_{2}+1\\ x_{1}+1\end{bmatrix}.$
The solution mapping $\Theta_{1}$ of the problem is as follows:
$\Theta_{1}(\xi_{1})=\left\{\begin{array}[]{cl}\left\{\Big{(}\frac{1}{1-2\xi_{1}},\frac{-1}{1-2\xi_{1}}\Big{)}\right\}&\ \ \hbox{ if }\ 0\leq\xi_{1}<\frac{1}{2}\ \hbox{ or }\ \frac{1}{2}<\xi_{1}\leq 1,\\ \emptyset&\ \ \hbox{ if }\qquad\qquad\ \,\xi_{1}=\frac{1}{2}.\\ \end{array}\right.$
Clearly, $\operatorname{dom}\Theta_{1}=[0,\frac{1}{2})\cup(\frac{1}{2},1]=\mathcal{O}$ is not connected but open in $[0,1]$. One observers that $\Theta_{1}$ is single-valued and continuous on $\mathcal{O}$. Its image has two connected components, both of them are unbounded.
Proposition 1
Assume that $\Psi$ is upper semi-continuous on $\operatorname{dom}\Psi$ and that $\Psi^{-1}$ also has connected values. Then, $\Psi(A)$ is disconnected if and only if $\operatorname{dom}\Psi$ is so.
Proof
Clearly, $\Psi^{-1}$ has closed graph and its image, namely $\operatorname{dom}\Psi\subseteq A$, is bounded. Thanks to Lemma 1, $\Psi^{-1}$ is upper semi-continuous on $\Psi(A)$. Hence, both $\Psi$ and $\Psi^{-1}$ are upper semi-continuous on their domains and have connected values. Therefore, the conclusion is implied from Theorem 2.1.∎
Example 4
We consider the mapping given in Example 3. The mapping satisfies the condition in Proposition 1. Clearly, the number of connected components of $\operatorname{dom}\Theta_{1}$ equals the one of the image $\Theta_{1}([0,1])$.
5 Application to vector optimization without generalized convexity
We study unboundedness of solution sets in vector optimization through solution mappings associated to specific scalarization methods. Let us recall solution concepts and fundamental results in vector optimization.
Let $K$ be a nonempty, closed and convex subset of $\operatorname{\mathbb{R}}^{n}$ and $f=(f_{1},\dots,f_{m})$ be a continuous mapping from $K$ to $\operatorname{\mathbb{R}}^{m}$. The vector minimization problem with the feasible set $K$, vector objective function $f$ and nonnegative orthant +$\operatorname{\mathbb{R}}^{m}_{+}$ as ordering cone is written formally as follows:
$\operatorname{VOP}(K,f)\qquad{\rm Minimize}\ \;f(x)\ \ {\rm subject\ to}\ \ x\in K.$
A point $x\in K$ is said to be a Pareto solution of $\operatorname{VOP}(K,f)$ if there exists no $y\in K$ such that $f(y)-f(x)$ belongs to +$-\operatorname{\mathbb{R}}^{m}_{+}\setminus\{0\}$. It is said to be a weak Pareto solution of $\operatorname{VOP}(K,f)$ if there does not exist a $y\in K$ such that $f(y)-f(x)$ belongs to $-\operatorname{int}\operatorname{\mathbb{R}}^{m}_{+}$. The Pareto solution set and weak Pareto solution set of this problem are respectively denoted by $\operatorname{Sol}(K,f)$ and $\operatorname{Sol^{w}}(K,f)$.
Remark 2
Let $x$ be given in $K$. It is $x\in\operatorname{Sol^{w}}(K,f)$ if and only if for any $y\in K$ there exists $k\in\{1,\dots,m\}$ such that $f_{k}(y)\geq f_{k}(x)$. Indeed, $x\in\operatorname{Sol^{w}}(K,f)$ is equivalent to the condition that for every $y\in K$ it is $f(y)-f(x)\not\in-\operatorname{int}\operatorname{\mathbb{R}}^{m}_{+}$, which is the same as $f_{k}(y)\geq f_{k}(x)$ for some index $k$.
If $f$ is a scalar function, so $f:K\to\operatorname{\mathbb{R}}$, than we point out the fact, that in this case the problem $\operatorname{VOP}(K,f)$ is the well known scalar minimization problem, by writing it as $\operatorname{OP}(K,f)$. In this case the (weak) Pareto solutions are the minimizers of $f$ over $K$.
Let $g:K\to\operatorname{\mathbb{R}}$ be a function. The function $g$ is said to be bounded from below by $\gamma\in\operatorname{\mathbb{R}}$ on $K$ if $g(x)\geq\gamma$ for all $x\in K$ and strictly bounded from below by $\gamma$ on $K$ if $g(x)>\gamma$ for all $x\in K$.
5.1 Two solution mappings
We now consider the $(m-1)$-standard simplex in $\operatorname{\mathbb{R}}^{m}$
$\Delta\coloneqq\{(\xi_{1},\dots,\xi_{m})\in\operatorname{\mathbb{R}}_{+}^{m}:\xi_{1}+\dots+\xi_{m}=1\}.$
Then its relative interior $\operatorname{ri}\Delta$ is the set of all $\xi$ in $\Delta$ whose coordinates are positive. Clearly, $\Delta$ is nonempty, convex and compact. For each $\xi$ in $\Delta$, we define
$f_{\xi}\coloneqq\xi_{1}f_{1}+\dots+\xi_{m}f_{m}.$
The solution mapping associated to this scalarization method of the vector optimization problem $\operatorname{VOP}(K,f)$ is defined and denoted as follows:
$\Theta_{1}(\xi)\coloneqq\operatorname{Sol}(K,f_{\xi}),$ (5.1)
where $\operatorname{Sol}(K,f_{\xi})$ is the solution set of the scalar problem $\operatorname{OP}(K,f_{\xi})$.
The mapping $\Theta_{1}$ is an effective tool to investigate the weak Pareto solution set and Pareto solution set of a vector optimization problem because of the following results which can be found in, for example, (ehrgott2005multicriteria, , Proposition 3.9) or (ansari2018vector, , Corollaries 4.1–4.3):
Lemma 4
The following inclusions hold for the vector optimization problem $\operatorname{VOP}(K,f)$:
[…]
Remark 3
The solution mapping $\Theta_{1}$ has closed graph. Indeed, consider a convergent sequence $(\xi^{k},x^{k})$ in
$\operatorname{gph}\Theta_{1}=\{(\xi,x)\in\Delta\times K:x\in\Theta_{1}(\xi)\}$
such that $(\xi^{k},x^{k})\to(\bar{\xi},\bar{x})$. By the closedness of $\Delta$ and $K$, $\bar{\xi}$ and $\bar{x}$ belong to $\Delta$ and $K$, respectively. Let $x\in K$ be arbitrarily given. For each $k$, one has $f_{\xi^{k}}(x)\geq f_{\xi^{k}}(x^{k})$. Take $k\to+\infty$, by the continuity of $f$, we get $f_{\bar{\xi}}(x)\geq f_{\bar{\xi}}(\bar{x})$. It follows that $\bar{x}\in\operatorname{Sol}(K,f_{\bar{\xi}})$.
We consider another solution mapping, which was introduced in warburton1983quasiconcave and is also very useful for investigating solution sets of vector optimization problems governed by quasi-convex functions. For $\xi$ in $\Delta$, we define
$f^{\xi}(x)\coloneqq\max\left\{\xi_{1}f_{1}(x),\ \dots\,,\ \xi_{m}f_{m}(x)\right\}.$ (5.3)
The solution mapping $\Theta_{2}:\Delta\rightrightarrows\operatorname{\mathbb{R}}^{n}$ associated to this scalarization is given by
$\Theta_{2}(\xi)=\operatorname{Sol}(K,f^{\xi}).$ (5.4)
Remark 4
One has $\Theta_{2}(\operatorname{ri}\Delta)\subseteq\operatorname{Sol^{w}}(K,f)$. Indeed, for any $x\in\Theta_{2}(\operatorname{ri}\Delta)$, there exists $\xi\in\operatorname{ri}\Delta$ with $x\in\operatorname{Sol}(K,f^{\xi})$. This implies that for every $y\in K$ there exists an index $k\in\{1,\dots,m\}$ such that for all $i\in\{1,\dots,m\}$ we have
$f^{\xi}(y)=\xi_{k}f_{k}(y)\geq\xi_{i}f_{i}(x).$
For $i=k$ we therefore get $\xi_{k}f_{k}(y)\geq\xi_{k}f_{k}(x)$. Since $\xi\in\operatorname{ri}\Delta$ we have $\xi_{k}>0$, so that we arrive at $f_{k}(y)\geq f_{k}(x)$, which shows $x\in\operatorname{Sol^{w}}(K,f)$ (see Remark 2).
Suppose that all objective functions $f_{1},\dots,f_{m}$ are even strictly bounded from below by 00 on $K$. Then, according to (warburton1983quasiconcave, , Theorem 3.4), we have the equality $\Theta_{2}(\operatorname{ri}\Delta)=\operatorname{Sol^{w}}(K,f)$. If additionally $\Theta_{2}$ is single-valued on its domain, then $\operatorname{Sol}(K,f)=\operatorname{Sol^{w}}(K,f)$, see (warburton1983quasiconcave, , Theorem 3.5).
In the following theorem, we prove that, if the objective functions are bounded from below by 00 on $K$, then the weak Pareto solution set is the image of $\Delta$ under $\Theta_{2}$. This theorem which is an improvement of (warburton1983quasiconcave, , Theorem 3.4) plays a vital role in the proof of Theorem 6.1 in the next section.
Theorem 5.1
If all objective functions $f_{1},\dots,f_{m}$ are bounded from below by 00 on $K$, then we have
$\operatorname{Sol^{w}}(K,f)=\Theta_{2}(\Delta).$ (5.5)
Proof
For $m=1$ this is evident, so from now on we assume $m>1$. To prove (5.5), we first prove $\operatorname{Sol^{w}}(K,f)\subseteq\Theta_{2}(\Delta)$. For this, we consider any $x\in\operatorname{Sol^{w}}(K,f)$. If there exists an index $k\in\{1,\dots,m\}$ with $f_{k}(x)=0$ then $x\in\operatorname{Sol}(K,f^{\xi})$, where $\xi=e^{k}$ is the $k$-th standard unit vector of $\operatorname{\mathbb{R}}^{m}$. Indeed, for any $y\in K$ one has $f^{\xi}(y)=\max\{0,f_{k}(y)\}\geq 0$. It is clear that $f^{\xi}(x)=0$, this implies $x\in\operatorname{Sol}(K,f^{\xi})$. From now on we assume $f(x)\in\operatorname{int}\operatorname{\mathbb{R}}^{m}_{+}$ and define $\xi\in\operatorname{\mathbb{R}}^{m}$ by
$\xi_{i}:=\frac{1/f_{i}(x)}{1/f_{1}(x)+\dots+1/f_{m}(x)},\quad i=1,\dots,m.$ (5.6)
Then it is $\xi\in\operatorname{ri}\Delta$ and we get $f^{\xi}(x)=\frac{1}{1/f_{1}(x)+\dots+1/f_{m}(x)}$. One has $\xi_{i}f_{i}(x)=\frac{1}{1/f_{1}(x)+\dots+1/f_{m}(x)}$, for all $i\in\{1,\dots,m\}$. Because of $x\in\operatorname{Sol^{w}}(K,f)$, for any $y\in K$, there exists an index $k\in\{1,\dots,m\}$ with $f_{k}(y)\geq f_{k}(x)$ implicating
$f^{\xi}(y)\geq\xi_{k}f_{k}(y)$ $=\frac{f_{k}(y)/f_{k}(x)}{1/f_{1}(x)+\dots+1/f_{m}(x)}$
$\geq\frac{1}{1/f_{1}(x)+\dots+1/f_{m}(x)}=f^{\xi}(x),$
which shows $x\in\operatorname{Sol}(K,f^{\xi})$.
We now prove $\Theta_{2}(\Delta)\subseteq\operatorname{Sol^{w}}(K,f)$. Remark 4 points out that $\Theta_{2}(\operatorname{ri}\Delta)\subseteq\operatorname{Sol^{w}}(K,f)$. Thus, we need only to show $\Theta_{2}(\partial_{r}\Delta)\subseteq\operatorname{Sol^{w}}(K,f)$. Consider $x\in\Theta_{2}(\partial_{r}\Delta)$. Again, there exists $\xi\in\partial_{r}\Delta$ such that $x\in\operatorname{Sol}(K,f^{\xi})$. If there exists an index $k\in\{1,\dots,m\}$ with $f_{k}(x)=0$ then because of the lower boundedness we immediately get $x\in\operatorname{Sol^{w}}(K,f)$, so that from now on we can assume $f(x)\in\operatorname{int}\operatorname{\mathbb{R}}^{m}_{+}$. Define
$I(\xi)\coloneqq\{i\in\{1,\dots,m\}\mid\xi_{i}>0\}.$
Then, since for all $i\in\{1,\dots,m\}\setminus I(\xi)$, we have $\xi_{i}f_{i}(x)=0$ it is
$f^{\xi}(x)=\max_{i\in I(\xi)}\xi_{i}f_{i}(x)>0.$ (5.7)
Now, as before, for every $y\in K$ there exists an index $k\in\{1,\dots,m\}$ such that for all $i\in\{1,\dots,m\}$ it is $\xi_{k}f_{k}(y)\geq\xi_{i}f_{i}(x)$ and from $f^{\xi}(x)>0$ we actually get $k\in I(\xi)$, which again allows to conclude $f_{k}(y)\geq f_{k}(x)$ and therefore, according to Remark 2, $x\in\operatorname{Sol^{w}}(K,f)$, which proves (5.5). ∎
Remark 5
In (warburton1983quasiconcave, , p. 543) it is stated that the claim of (warburton1983quasiconcave, , Theorem 3.4) remains correct for $f$ bounded from below by any value $\gamma$. However, we want to point out that the bound zero is of utmost importance and cannot be generalized without further assumptions. In the following, we can see a counter-example.
Example 5
Consider $f:\operatorname{\mathbb{R}}^{2}\to\operatorname{\mathbb{R}}^{2}$, $f(x)\coloneqq x$ and $K\coloneqq[-1,0]\times[0,1]$. It can easily be seen that
$\operatorname{Sol^{w}}(K,f)=\big{(}[-1,0]\times\{0\}\big{)}\cup\big{(}\{-1\}\times[0,1]\big{)}.$
For any $\xi\in\Delta$ and $x\in K$, it is $\max\{\xi_{1}x_{1},\xi_{2}x_{2}\}=\xi_{2}x_{2}$, because the first part of the maximum is always non-positive while the second part is always non-negative. Therefore, we have
$\Theta_{2}(\xi)=\begin{cases}[-1,0]\times\{0\}&\text{if}\ \xi_{2}\neq 0\\ K&\text{if}\ \xi_{2}=0\end{cases}$
which gives
$\Theta_{2}(\operatorname{ri}\Delta)=[-1,0]\times\{0\}\subsetneq\operatorname{Sol^{w}}(K,f)\subsetneq K=\Theta_{2}(\partial_{r}\Delta)=\Theta_{2}(\Delta).$
To illustrate the sets, in Figure 3, $\Theta_{2}(\operatorname{ri}\Delta)$ is colored by blue and dashed, and $\operatorname{Sol^{w}}(K,f)$ is colored by red.
[…]
Figure 3: The relationship $\Theta_{2}(\operatorname{ri}\Delta)\varsubsetneq\operatorname{Sol^{w}}(K,f)\varsubsetneq\Theta_{2}(\operatorname{ri}\Delta)$.
Remark 6
Similarly to $\Theta_{1}$, the solution mapping $\Theta_{2}$ also has closed graph. To demonstrate this remark, we consider a convergent sequence $(\xi^{k},x^{k})$ in
$\operatorname{gph}\Theta_{2}=\{(\xi,x)\in\Delta\times K:x\in\Theta_{2}(\xi)\}$
such that $(\xi^{k},x^{k})\to(\bar{\xi},\bar{x})$. By the closedness of $\Delta$ and $K$, $\bar{\xi}$ and $\bar{x}$ belong to $\Delta$ and $K$, respectively. Let $x\in K$ be arbitrarily given. For each $k$, one has
$\max\left\{\xi^{k}_{1}f_{1}(x),\ \dots,\ \xi^{k}_{m}f_{m}(x)\right\}\ \geq\ \max\left\{\xi^{k}_{1}f_{1}(x^{k}),\ \dots,\ \xi^{k}_{m}f_{m}(x^{k})\right\}.$
Take $k\to+\infty$. By the continuity of $f$, we get $f^{\bar{\xi}}(x)\geq f^{\bar{\xi}}(\bar{x})$. It follows that $\bar{x}\in\operatorname{Sol}(K,f^{\bar{\xi}})$.
5.2 Problems without imposing any generalized convexity
We emphasize that, in the following theorem, we do not need any generalized convexity assumptions for the objective functions. By Remarks 3 and 6 the results obtained in Section 3 can be applied to both $\Theta_{1}$ and $\Theta_{2}$. We prefer to omit a detailed formulation of the results related to $\Theta_{2}$, which requires additionally that the objective functions are bounded from below by 00.
Theorem 5.2
Assume that the domain of the solution mapping $\Theta_{1}$ given in (5.1) is non-closed. Then, the weak Pareto solution set is unbounded. Moreover, the Pareto solution set is unbounded provided that $\Theta_{1}(\operatorname{ri}\Delta)$ is nonempty.
Proof
According to Remark 3, the solution mapping $\Theta_{1}$ has closed graph. The conclusions are obtained straightforwardly by applying the conclusions in Theorem 3.1 and inclusions in (5.2). ∎
Example 6
Consider the bicriteria optimization problem $\operatorname{VOP}(K,f)$ in $\operatorname{\mathbb{R}}^{2}$, with the feasible $K=\operatorname{\mathbb{R}}^{2}$ and two objective functions given by
$f_{1}(x)=x_{1}^{4}-2x_{2},\quad f_{2}(x)=-2x_{1}^{2}+x_{2}^{2}.$
One can see that the functions are not bounded from below on $K$. Thus, two initial points $(0,1),(1,0)$ in $\Delta$ do not belong to the domain of the solution mapping $\Theta_{1}$ of $\operatorname{VOP}(K,f)$. Let $\xi\in\operatorname{ri}\Delta$, equivalently $\xi_{1}\in(0,1)$. One has $f_{\xi}=\xi_{1}f_{1}+(1-\xi_{1})f_{2}$ and an easy computation shows that
$f_{\xi}(x_{1},x_{2})=\xi_{1}\left(x_{1}^{2}-\frac{1-\xi_{1}}{\xi_{1}}\right)^{2}+(1-\xi_{1})\left(x_{2}-\frac{\xi_{1}}{1-\xi_{1}}\right)^{2}+h(\xi_{1}),$
where $h$ is a function depended only on $\xi_{1}$. Since the first two terms in $f_{\xi}$ are nonnegative, $f_{\xi}$ reaches its minimum if and only if these terms are zero. Hence, the solution mapping of $\operatorname{VOP}(K,f)$ is as given in (3.1) in Example 1 and its domain is
$\operatorname{dom}\Theta_{1}=\Delta\setminus\{(0,1),(1,0)\}=\operatorname{ri}\Delta.$
This set is non-closed and its image is given as in (3.2). Hence, according to Theorem 5.2, the Pareto solution set and weak Pareto solution set of the problem are unbounded as $\operatorname{Sol}(K,f)\supseteq\Theta_{1}([0,1])$.
Remark 7
Let $\operatorname{\mathcal{A}}$ be a connected component of $\Theta_{1}(\Delta)$, where $\Theta_{1}$ is given in (5.1). If $\Theta_{1}^{-1}(\operatorname{\mathcal{A}})$ is non-closed, then $\operatorname{\mathcal{A}}$ is unbounded. Indeed, according to Remark 3, $\Theta_{1}$ has closed graph. By applying straightforwardly Remark 1, we obtain the conclusion.
6 Application to quasi-convex vector optimization
We now recall some notions related to generalized convexity. We refer to the monographs ansari2013generalized ; avriel2010generalized for a general introduction to generalized convexity.
Consider a function $g:K\to\operatorname{\mathbb{R}}$. This function is quasi-convex on $K$ if for $x,y$ in $K$, $x\neq y$, it is
$g(tx+(1-t)y)\leq\max\{g(x),g(y)\},$ (6.1)
for all $t\in(0,1)$. Equivalently, $g$ is quasi-convex if and only if the level set $[g\leq\lambda]$ is a convex set for all $\lambda\in\operatorname{\mathbb{R}}$. Hence, the solution set of the optimization problem $\operatorname{OP}(K,g)$ is convex. In addition, when the inequality in (6.1) is replaced by strict one, we say that $g$ is strictly quasi-convex on $K$. Recall that if $g$ is strictly quasi-convex then the solution set of $\operatorname{OP}(K,g)$ has at most one element.
6.1 Stability of solutions in parametric optimization under quasi-convexity
Let $T$ be a convex subset of $\operatorname{\mathbb{R}}^{k}$ and $g:K\times T\to\operatorname{\mathbb{R}}$ be continuous on $K\times T$. We consider the parametric optimization problem $\operatorname{OP}(K,g(\cdot,t))$, where $t\in T$, and define the solution mapping $\Theta$ as follows:
$\Theta(t)\coloneqq\operatorname{Sol}(K,g(\cdot,t)).$ (6.2)
The following lemma about stability of solutions in quasi-convex optimization plays an important role in the proof of the main result, see Theorem 6.1; Its proof can be found in (bank1983non, , Theorem 4.3.3):
Lemma 5
Assume that, for any parameter $t\in T$, $g(\cdot,t)$ is quasi-convex on $K$. Then, the solution mapping $\Theta$ is upper semicontinuous at all points $t\in T$ such that $\Theta(t)$ is nonempty and bounded.
Considering the above parametric problem under strict quasi-convexity condition, we have the following lemma which will be used in the proof of Corollary 1.
Lemma 6
Assume that, for any parameter $t\in T$, $g(\cdot,t)$ is strictly quasi-convex on $K$. Then, the solution map $\Theta$ is single-valued and continuous on its domain which is open in $T$.
Proof
The single-valuedness of $\Theta$ is implied from the strictly quasi-convexity of $g(\cdot,t)$ on $K$. Hence, for arbitrary $t\in\operatorname{dom}\Theta$, $\Theta(t)$ is the unique isolated local solution of $\operatorname{OP}(K,g(\cdot,t))$. Applying (klatte1985stability, , Theorem 1) to the considered parametric optimization problem, we conclude that there exists a bounded and open set $V\supseteq\Theta(t)$ and a neighborhood $U$ of $t$ such that, for all $t^{\prime}\in U$, $\Theta(t^{\prime})$ is nonempty and contained in $V$. Hence, the domain of $\Theta$ is open and $\Theta$ is continuous on this set. ∎
6.2 Problems governed by quasi-convex functions
Under quasi-convexity, to ensure disconnectedness of all components of the weak Pareto solution set, the authors in (hoa2007unbounded, , Theorem 3.2) used a condition which says that the number of components of the set is finite. In the following theorem, we propose two other conditions concerning the solution mapping $\Theta_{2}$ given in (5.4) to replace it. Let
$\mathcal{O}=\{\xi\in\operatorname{dom}\Theta_{2}:\Theta_{2}(\xi)\text{ is bounded}\}.$ (6.3)
Theorem 6.1
Let $f_{1},\dots,f_{m}$ be quasi-convex and bounded from below by 00 on $K$. Assume that one of the following conditions holds:
$\operatorname{dom}\Theta_{2}$ is connected;
$\mathcal{O}$ is open in $\Delta$.
If $\operatorname{Sol^{w}}(K,f)$ is disconnected, then all the components are unbounded. Consequently, if the set is bounded then it is connected.
Proof
Since the objective functions are bounded from below by 00 on $K$, thanks to Theorem 5.1, $\operatorname{Sol^{w}}(K,f)=\Theta_{2}(\Delta)$. Moreover, $f_{1},\dots,f_{m}$ are quasi-convex, from (ansari2013generalized, , Theorem 1.31), $f^{\xi}$ given in (5.3) is also quasi-convex on $K$, for $\xi\in\Delta$, and then $\Theta_{2}$ has convex values.
According to Lemma 5, $\Theta_{2}$ is upper semi-continuous on $\mathcal{O}$. Therefore, the desired conclusion is obtained by applying directly Theorems 4.1 and 4.2 for $\Theta_{2}$ with the note that $\Theta_{2}$ has closed graph and connected values. ∎
If the solution mapping $\Theta_{2}$ is single-valued on its domain, we have the following result:
Corollary 1
Let all $f_{i}$ be quasi-convex and strictly bounded from below by 00 on $K$. Suppose that $\Theta_{2}$ is single-valued on its domain. If $\operatorname{Sol}(K,f)$ is disconnected then all its components are unbounded.
Proof
From the assumption, the domain of $\Theta_{2}$ is $\mathcal{O}$, where $\mathcal{O}$ is defined by (6.3). Hence, $\mathcal{O}$ is the set of $\xi$ in $\Delta$ such that $\Theta_{2}(\xi)$ has only one element. Thanks to (warburton1983quasiconcave, , Theorem 3.5), we claim that $\operatorname{Sol}(K,f)$ coincides with $\operatorname{Sol^{w}}(K,f)$.
By applying Lemma 6 for the problem $\operatorname{VOP}(K,f)$, we conclude that $\operatorname{dom}\Theta_{2}$ is open in $\Delta$. As $\operatorname{VOP}(K,f)$ satisfies condition $\rm(b)$ in Theorem 6.1, by applying that theorem for $\operatorname{Sol^{w}}(K,f)$, we obtain the desired conclusion. ∎
Example 7
We consider the problem given in (warburton1983quasiconcave, , Example 5.1) whose constraint set is
$K=\{x\in\operatorname{\mathbb{R}}^{2}:2\leq x_{1},0\leq x_{2}\leq 4\}$
and objective functions are
$f_{1}(x)=\frac{x_{1}}{x_{1}+x_{2}-1},\ \ f_{2}(x)=\frac{x_{1}}{x_{1}-x_{2}+3}.$
Both $f_{1},f_{2}$ are strictly bounded from below by 00 on $K$. This is a quasi-convex problem and its solution mapping $\Theta_{2}$ is given by
$\Theta_{2}(\xi_{1})=\left\{\begin{array}[]{cl}\emptyset&\ \ \hbox{ if }\ 0\leq\xi_{1}<\frac{1}{2}\hbox{ or }\ \frac{5}{6}<\xi_{1}\leq 1,\\ \left\{\Big{(}\frac{1+2\xi_{1}}{1-2\xi_{1}},0\Big{)}\right\}&\ \ \hbox{ if }\qquad\quad\ \frac{1}{6}\leq\xi_{1}<\frac{1}{2},\\ \left\{\Big{(}\frac{3-2\xi_{1}}{2\xi_{1}-1},4\Big{)}\right\}&\ \ \hbox{ if }\qquad\quad\ \frac{1}{2}<\xi_{1}\leq\frac{5}{6}.\end{array}\right.$
Clearly, the domain of $\Theta_{2}$ is the union of two open intervals $\left(\frac{1}{6},\frac{1}{2}\right)\cup\left(\frac{1}{2},\frac{5}{6}\right)$; it is disconnected. The mapping is single-valued on its domain. According to Corollary 1, the Pareto solution set and weak Pareto solution set are coincident. They are the union of two disconnected parallel lines:
$\{(x_{1},x_{2})\in\operatorname{\mathbb{R}}^{2}:x_{1}\geq 2,x_{2}=0\}\cup\{(x_{1},x_{2})\in\operatorname{\mathbb{R}}^{2}:x_{1}\geq 2,x_{2}=4\}.$
Remark 8
If all $f_{i}$ are convex, we have the same conclusions in Theorem 6.1 for the solution map $\Theta_{1}$ defined by (5.1). However, until now, we have not known any explicit convex vector optimization problems whose (weak) Pareto solution set is disconnected. This is related to Question 9.8 in yen2011monotone and Question 5.1 in hieu2021disconnectedness .
In the case that all objective functions are strictly quasi-convex, $\Theta_{2}$ is single-valued on its domain. Therefore, we have the following theorem:
Theorem 6.2
Suppose that all $f_{i}$ are strictly quasi-convex and strictly bounded from below by 00 on $K$. Then, one has $\operatorname{Sol}(K,f)=\operatorname{Sol^{w}}(K,f)$. Moreover, if $\operatorname{Sol}(K,f)$ is disconnected then all its components are unbounded.
Proof
For $\xi\in\Delta$, $f^{\xi}$ is also strictly quasi-convex on $K$ (see (warburton1983quasiconcave, , Lemma 4.2)); Hence $\Theta_{2}$ is single-valued on its domain. The conclusions are implied straightforwardly from Corollary 1. ∎
Perspectives and future work
The results related to set-valued mappings having connected values in this paper can be useful for investigating (un)bounedness and (un)connectedness of solution sets in set optimization, vector equilibrium, or vector variational inequality under some generalized convexity/monotone assumptions. We plan to exploit special structures such as polyhedral or semi-algebraic to discover new properties concerning contractibility or connectedness of solution sets of these problems as in huong2017connectedness ; huong2016polynomial ; hieu2019polynomialVVIS ; hieu2020application .

### v2

@@
-7491, Trondheim, Norway
+7491, Trondheim, Norway Center for Advanced Intelligence Project, RIKEN 103-0027, Tokyo, Japan
 Abstract
@@
 1 Introduction
-Studying structure of solutions plays an important role in vector optimization theory, see, for example (lucbook, , Chapter 6) and (ehrgott2005multicriteria, , Chapter 3).
-The most effective method for such studying is considering solution sets as the images of solution mappings based on scalarization techniques.
+Studying the structure of solution sets plays an important role in vector optimization theory; see, for example (lucbook, , Chapter 6) and (ehrgott2005multicriteria, , Chapter 3).
+The most effective method for such studies is to consider the solution sets in question as the images of the solution mappings obtained by some scalarization techniques.
 It is well-known that the image of a connected set under a continuous single-valued mapping is connected.
-Naccache in naccache1978connectedness , Warburton in warburton1983quasiconcave , and Hiriart-Urruty in hiriart1985images extended the result for a set-valued mapping having connected values, provided that the mapping is upper (or lower) semi-continuous, see Theorem 2.1.
-This criterion opened a door to study connectedness of solution sets of not only vector optimization problems but also of set optimization problems, vector variational inequalities, vector equilibrium problems.
-There are numerous works related to this topic, we can only list a few of them, for example luc1987connectedness ; sun1996connectedness ; gong2007connectedness ; yen2011monotone ; han2018existence ; sharma2022connectedness ; peng2019connectedness ; han2022connectedness ; anh2022connectedness .
-Almost all of the obtained results require generalized convexity or monotonicity assumptions to ensure connected-valuedness, and require boundedness of the constraint sets or certain coercivity conditions to guarantee semi-continuity of solution mappings.
+Naccache in naccache1978connectedness , Warburton in warburton1983quasiconcave , and Hiriart-Urruty in hiriart1985images extended the result for a set-valued mapping having connected values, provided that the mapping is upper (or lower) semi-continuous; see Theorem 2.1.
+This criterion opened a door for studying connectedness of solution sets of not only vector optimization problems but also set optimization problems, vector variational inequalities and vector equilibrium problems.
+There are numerous works related to this topic, we can only list a few of them, for example luc1987connectedness ; sun1996connectedness ; gong2007connectedness ; yen2011monotone ; han2018existence ; sharma2023connectedness ; peng2019connectedness ; han2022connectedness ; anh2022connectedness .
+Almost all therein obtained results require generalized convexity or monotonicity assumptions to ensure connected-valuedness, and require boundedness of the constraint sets or certain coercivity conditions to guarantee semi-continuity of solution mappings.
 The aim of this paper is to look for tools that can be applied to investigate unboundedness and disconnectedness of solutions to problems governed by unbounded data in optimization theory.
@@
 We obtain new properties related to unboundedness of the (weak) Pareto solution set of a problem in both cases: with a generalized convexity assumption and without imposing generalized convexity.
-For the second case, in Theorem 5.2 we prove that if the domain of a solution mapping between two aforementioned ones is non-closed then the weak Pareto solution set is unbounded.
-For a problem governed by semi-strictly quasi-convex functions, to ensure disconnectedness of all components of the weak Pareto solution set, Hoa, Huy, Phuong and Yen in (hoa2007unbounded, , Theorem 3.2) used a condition which says that the number of components of the set is finite.
-In Theorem 6.1, we introduce two other criteria for quasi-convex problems.
-As a result, when the objective functions are strictly quasi-convex, the Pareto solution set and weak Pareto solution set are coincident, and if this set is disconnected then all the components are unbounded.
+For the second case, in Theorem 5.2 we prove that if the domain of one of the two aforementioned solution mappings (denoted by $\Theta_{1}$ and $\Theta_{2}$; see the definitions in Subsection 5.1) is non-closed then the weak Pareto solution set is unbounded.
+The reverse implication holds for $\Theta_{1}$ if the problem is convex and this solution map has bounded values, see Theorem 6.3.
+In Theorem 6.2, we have the same conclusion for $\Theta_{2}$ when the objective functions are quasi-convex and non-negative over the constraint set.
+For a problem governed by semi-strictly quasi-convex functions, to ensure disconnectedness of all components of the weak Pareto solution set, Hoa et al.
+(hoa2007unbounded, , Theorem 3.2) used the condition that the number of components of the set is finite.
+In Theorem 6.1, we introduce two other criteria for this property of quasi-convex problems.
+As a result, when the objective functions are strictly quasi-convex, the Pareto solution set coincides with the weak Pareto solution set, and if this set is disconnected then all the components are unbounded.
 The remaining part of this paper consists of five sections.
@@
 Let $X$ be a nonempty subset of $\operatorname{\mathbb{R}}^{m}$.
-We consider set-valued mapping $S$ from $X$ to $\operatorname{\mathbb{R}}^{n}$, $S:X\rightrightarrows\operatorname{\mathbb{R}}^{n}$.
+We consider a set-valued mapping $S$ from $X$ to $\operatorname{\mathbb{R}}^{n}$, $S:X\rightrightarrows\operatorname{\mathbb{R}}^{n}$.
 The domain of $S$ is the set of all elements $u$ in $X$ such that $S(u)$ is non-empty:
@@
 $\operatorname{gph}S\coloneqq\{(u,v)\in X\times\operatorname{\mathbb{R}}^{n}:v\in S(u)\}$
-is a closed subset of $X\times\operatorname{\mathbb{R}}^{n}$ then one says that $S$ has closed graph.
-If $S$ has closed graph then $S(u)$ is a closed set for all $u\in X$.
+is a closed subset of $X\times\operatorname{\mathbb{R}}^{n}$, then one says that $S$ has closed graph.
+If $S$ has closed graph, then $S(u)$ is a closed set for every $u\in X$.
 The inverse mapping of $S$, denoted by $S^{-1}:\operatorname{\mathbb{R}}^{n}\rightrightarrows X$, is defined by
@@
 The mapping $S$ is locally bounded at $\bar{u}$ if there exists an open neighborhood $U$ of $\bar{u}$ such that $S(U\cap X)$ is bounded.
-If $S$ is locally bounded at every $u\in X$ then $S$ is said to be locally bounded on $X$.
+If $S$ is locally bounded at every $u\in X$, then $S$ is said to be locally bounded on $X$.
 The mapping $S$ is upper semi-continuous at $\bar{u}\in X$ if for any open set $V\subseteq\operatorname{\mathbb{R}}^{n}$ such that $S(\bar{u})\subseteq V$ there exists an open neighborhood $U$ of $\bar{u}$ such that $S(u)\subseteq V$ for all $u\in U$.
-If $S$ is upper semi-continuous at every $u\in X$ then $S$ is said to be upper semi-continuous on $X$.
+If $S$ is upper semi-continuous at every $u\in X$, then $S$ is said to be upper semi-continuous on $X$.
 From (AuslenderTeboulle2003AsymptoticConesFunctionsOptimizationVariationalInequalitiesbook, , Theorem 1.4.1), one has the following lemma which will be used in the proof of Theorem 4.1.
@@
 Let $Y$ be a nonempty subset of $\operatorname{\mathbb{R}}^{n}$.
-One says that $Y$ is connected if there do not exist two nonempty disjoint subsets $A,B$ of $Y$ with $Y\subseteq A\cup B$ and two open subsets $U,V$ in $\operatorname{\mathbb{R}}^{n}$ such that $A\subseteq U$, $B\subseteq V$ and $U\cap V=\emptyset$.
+One says that $Y$ is connected if there do not exist two nonempty disjoint subsets $A,B$ of $Y$ with $Y\subseteq A\cup B$ and two open subsets $U,V$ in $\operatorname{\mathbb{R}}^{n}$ such that $A\subseteq U$, $B\subseteq V$ and $(U\cap Y)\cap(V\cap Y)=\emptyset$.
 Otherwise $Y$ is called disconnected.
 A nonempty subset $A\subseteq Y$ is said to be a connected component of $Y$ if $A$ is connected and it is not a proper subset of any connected subset of $Y$.
-Recall that when $Y$ is connected, if the set $A\subseteq Y$ is closed and open in $Y$ then $A=Y$.
+Recall that when $Y$ is connected, if the set $A\subseteq Y$ is closed and open in $Y$, then $A=Y$.
 Moreover, if $Y$ is connected then the union $Y\cup\{y\}$, where $y$ belongs to the closure $\operatorname{cl}(Y)$, is also connected.
-One says that the set-valued mapping $S$ has connected (convex, bounded) values if $S(x)$ is connected (convex, bounded) for every $x$ on the domain of $S$.
-In order to be convenient for readers to compare our developments with the before-mentioned result by Naccache, Warburton, and Hiriart-Urruty, we recall the following theorem; its proofs can be found in (naccache1978connectedness, , Lemma 2.1), (warburton1983quasiconcave, , Theorem 3.1), or (hiriart1985images, , Theorem 3.1):
+One says that the set-valued mapping $S$ has connected (convex, bounded) values if $S(u)$ is connected (convex, bounded) for every $u$ in the domain of $S$.
+For the convenience of the readers in comparing our developments with the before-mentioned result by Naccache, Warburton, and Hiriart-Urruty, we recall the following theorem; its proofs can be found in (naccache1978connectedness, , Lemma 2.1), (warburton1983quasiconcave, , Theorem 3.1), or (hiriart1985images, , Theorem 3.1).
 Theorem 2.1
-Assume that $X$ is connected and $S$ has connected values. If $S$ is upper semi-continuous on $X$, then the image $S(X)$ is connected.
+Assume that $\operatorname{dom}S$ is connected and $S$ has connected values. If $S$ is upper semi-continuous on $X$, then the image $S(X)$ is connected.
 3 Set-valued mappings having non-closed domains
-Let $A$ be a nonempty subset of $\operatorname{\mathbb{R}}^{m}$. We want to emphasize that, throughout the present section, the set-valued mapping
-$\Phi:A\rightrightarrows\operatorname{\mathbb{R}}^{n},\ a\mapsto\Phi(a),$
-has closed graph. We denote by $\operatorname{ri}A$ its relative interior.
-The proofs of the main results in the present and next sections are based on the following lemma:
+Let $A$ be a nonempty subset of $\operatorname{\mathbb{R}}^{m}$.
+We denote by $\operatorname{ri}A$ its relative interior.
+Throughout the present section, it is assumed that the following set-valued mapping has closed graph:
+$\Phi:A\rightrightarrows\operatorname{\mathbb{R}}^{n},\ a\mapsto\Phi(a).$
+The proofs of the main results in the present and next sections are based on the following lemma.
 Lemma 2
 If $\operatorname{\mathcal{A}}$ is a bounded connected component of $\Phi(A)$, then $\Phi^{-1}(\operatorname{\mathcal{A}})$ is closed.
-Furthermore, if $\Phi(A)$ is a bounded, then $\operatorname{dom}\Phi$ is closed.
+Furthermore, if $\Phi(A)$ is bounded, then $\operatorname{dom}\Phi$ is closed.
 Proof
 Let $\operatorname{\mathcal{A}}$ be a bounded connected component of $\Phi(A)$.
-To prove that $\Phi^{-1}(\operatorname{\mathcal{A}})$ is closed, we suppose that the sequence $\{a^{k}\}\subseteq\Phi^{-1}(\operatorname{\mathcal{A}})$ converges to $\bar{a}\in\operatorname{\mathbb{R}}^{m}$ and need to prove $\bar{a}\in\Phi^{-1}(\operatorname{\mathcal{A}})$.
-Indeed, let $\{x^{k}\}\subseteq\operatorname{\mathcal{A}}$ be a sequence such that $x^{k}\in\Phi(a^{k})$ for every $k\in\operatorname{\mathbb{N}}$.
+To prove that $\Phi^{-1}(\operatorname{\mathcal{A}})$ is closed, we suppose that a sequence $\{a^{k}\}\subseteq\Phi^{-1}(\operatorname{\mathcal{A}})$ converges to $\bar{a}\in\operatorname{\mathbb{R}}^{m}$ and need to prove $\bar{a}\in\Phi^{-1}(\operatorname{\mathcal{A}})$.
+Let $\{x^{k}\}\subseteq\operatorname{\mathcal{A}}$ be a sequence such that $x^{k}\in\Phi(a^{k})$ for every $k\in\operatorname{\mathbb{N}}$.
 By the boundedness of $\operatorname{\mathcal{A}}$, we can assume that $x^{k}\to\bar{x}$.
@@
 This means that $\bar{x}\in\operatorname{\mathcal{A}}$.
-We thus get $\bar{x}\in\operatorname{\mathcal{A}}\cap\Phi(\bar{a})\neq\emptyset$, and then $\bar{a}\in\Phi^{-1}(\operatorname{\mathcal{A}})$ by definition.
+We thus get $\bar{x}\in\operatorname{\mathcal{A}}\cap\Phi(\bar{a})$, and then $\bar{a}\in\Phi^{-1}(\operatorname{\mathcal{A}})$ by definition.
 Therefore, the inverse image $\Phi^{-1}(\operatorname{\mathcal{A}})$ is closed.
-The second conclusion is proven similarly. ∎
+We now prove the second conclusion.
+Suppose that $\Phi(A)$ is bounded.
+To prove closedness of $\operatorname{dom}\Phi$, we suppose that a sequence $\{a^{k}\}\subseteq\operatorname{dom}\Phi$ converges to $\bar{a}\in\operatorname{\mathbb{R}}^{m}$ and need to prove $\bar{a}\in\operatorname{dom}\Phi$.
+Let $\{x^{k}\}\subseteq\Phi(A)$ be a sequence such that $x^{k}\in\Phi(a^{k})$ for every $k\in\operatorname{\mathbb{N}}$.
+By the boundedness of $\Phi(A)$, we can assume that $x^{k}\to\bar{x}$.
+Since the graph of $\Phi$ is closed, $(\bar{a},\bar{x})$ belongs to $\operatorname{gph}\Phi$, thus $\bar{x}\in\Phi(\bar{a})$.
+This implies that $\bar{a}$ belongs to $\operatorname{dom}\Phi$.
+∎
 Theorem 3.1
@@
 Then, $\Phi(A)$ is unbounded.
-Moreover, if additionally $A$ is convex and $\Phi(\operatorname{ri}A)$ is nonempty then the latter set is also unbounded.
+Moreover, if additionally $(\operatorname{ri}A)\cap\operatorname{dom}\Phi$ is dense in $\operatorname{dom}\Phi$, then $\Phi(\operatorname{ri}A)$ is unbounded.
 Proof
 By the non-closedness of $\operatorname{dom}\Phi$ and the second conclusion in Lemma 2, we conclude straightforwardly that $\Phi(A)$ is unbounded. We only need to prove the second claim.
-Now let $A$ be convex. Suppose that $\Phi(\operatorname{ri}A)$ is nonempty but on the contrary the set is bounded. Let $\overline{\Sigma}$ be the closure of the following set:
+Let $(\operatorname{ri}A)\cap\operatorname{dom}\Phi$ be dense in $\operatorname{dom}\Phi$.
+Suppose on the contrary that $\Phi(\operatorname{ri}A)$ is bounded.
+Let $\overline{\Sigma}$ be the closure of the following set:
 $\Sigma\coloneqq\left(\operatorname{gph}\Phi\right)\cap\left(\operatorname{ri}A\times\operatorname{\mathbb{R}}^{n}\right)=\{(a,x)\in\operatorname{ri}A\times\operatorname{\mathbb{R}}^{n}:x\in\Phi(a)\}$.
@@
 Let $\pi:\operatorname{\mathbb{R}}^{m+n}\to\operatorname{\mathbb{R}}^{n}$ be the projection defined by $\pi(t_{1},\dots,t_{m},t_{m+1},\dots,t_{m+n})=(t_{m+1},\dots,t_{m+n})$.
-Clearly, the image of $\overline{\Sigma}$ under $\pi$ is bounded and
-$\Phi(\operatorname{ri}A)\subseteq\pi(\overline{\Sigma})\subseteq\Phi(A).$
-We consider a new mapping $\Phi^{\prime}$ defined by
+It is not difficult to see that
+$\Phi(\operatorname{ri}A)\subseteq\pi(\overline{\Sigma})\subseteq\overline{\pi(\Sigma)}.$
+Moreover, by the boundedness of $\Phi(\operatorname{ri}A)$ and the definition of $\Sigma$, we see that $\pi(\Sigma)$ is bounded, hence $\overline{\pi(\Sigma)}$ is as well.
+Thus, we claim that the image of $\overline{\Sigma}$ under $\pi$ is bounded.
+We now consider a new mapping $\Phi^{\prime}$ defined by
 $\Phi^{\prime}:A\rightrightarrows\operatorname{\mathbb{R}}^{n},\ \Phi^{\prime}(a)\coloneqq\pi(\overline{\Sigma})\cap\Phi(a).$
@@
 By boundedness of $\pi(\overline{\Sigma})$, $\Phi^{\prime}(A)=\pi(\overline{\Sigma})$ is also bounded.
-Applying the second conclusion in Lemma 2 to $\Phi^{\prime}$, $\operatorname{dom}\Phi^{\prime}$ is closed.
+Applying the second conclusion in Lemma 2 to $\Phi^{\prime}$, we get that $\operatorname{dom}\Phi^{\prime}$ is closed.
 Furthermore, it is not difficult to see the following inclusions:
 $(\operatorname{ri}A)\cap\operatorname{dom}\Phi\ \subseteq\ \operatorname{dom}\Phi^{\prime}\ \subseteq\ \operatorname{dom}\Phi.$
-By convexity of $A$ and density of $(\operatorname{ri}A)\cap\operatorname{dom}\Phi$ in $\operatorname{dom}\Phi$, $\operatorname{dom}\Phi^{\prime}$ is also dense in $\operatorname{dom}\Phi$.
+By the density of $(\operatorname{ri}A)\cap\operatorname{dom}\Phi$ in $\operatorname{dom}\Phi$, the above inclusions imply that $\operatorname{dom}\Phi^{\prime}$ is also dense in $\operatorname{dom}\Phi$.
 It follows that $\operatorname{dom}\Phi$ is closed.
@@
 Example 1
-From Example 6 we consider the set-valued mapping that is the solution mapping $\Theta_{1}:\Delta\rightrightarrows\operatorname{\mathbb{R}}^{2}$ of the there given vector optimization problem:
-$\Theta_{1}(\xi_{1})=\left\{\begin{array}[]{cl}\left\{\Big{(}\pm\sqrt{\frac{1-\xi_{1}}{\xi_{1}}},\frac{\xi_{1}}{1-\xi_{1}}\Big{)}\right\}&\ \ \hbox{ if }\ \xi_{1}\in(0,1),\\ \emptyset&\ \ \hbox{ if }\ \xi_{1}\in\{0,1\}.\\ \end{array}\right.$ (3.1)
+From Example 6 we consider the set-valued mapping that is the solution mapping $\Theta_{1}:\Delta\rightrightarrows\operatorname{\mathbb{R}}^{2}$ of the vector optimization problem given there:
+$\Theta_{1}(\xi_{1})=\left\{\begin{array}[]{cl}\left\{\Big{(}\pm\sqrt{\frac{1-\xi_{1}}{\xi_{1}}},\frac{\xi_{1}}{1-\xi_{1}}\Big{)}\right\}&\ \ \hbox{ if }\ \xi_{1}\in(0,1),\vskip 3.0pt plus 1.0pt minus 1.0pt\\ \emptyset&\ \ \hbox{ if }\ \xi_{1}\in\{0,1\}.\\ \end{array}\right.$ (3.1)
 We can see that the graph of $\Theta_{1}$ is closed and that the domain of $\Theta_{1}$ is the open interval $(0,1)$ which is non-closed.
@@
 In particular, we have
-$\Theta_{1}([0,1])=\big{\{}(x_{1},x_{2})\in\operatorname{\mathbb{R}}^{2}:x_{1}^{2}x_{2}=1\big{\}}=\operatorname{\mathcal{A}}\dot{\cup}\operatorname{\mathcal{B}},$ (3.2)
+$\Theta_{1}([0,1])=\big{\{}(x_{1},x_{2})\in\operatorname{\mathbb{R}}^{2}:x_{1}^{2}x_{2}=1\big{\}}=\operatorname{\mathcal{A}}\cup\operatorname{\mathcal{B}},$ (3.2)
 where
@@
 4 Set-valued mappings having connected values
-It is important to underline that, throughout the section, we consider another set-valued mapping
+Throughout the section, we consider another set-valued mapping
 $\Psi:A\rightrightarrows\operatorname{\mathbb{R}}^{n},\ a\mapsto\Psi(a),$
@@
 On the contrary, we assume that $\operatorname{\mathcal{A}}$ is bounded.
-Let $\operatorname{\mathcal{B}}$ is another connected component of $\Psi(A)$.
-Then there exists an open set $V$ containing $\operatorname{\mathcal{A}}$ such that $V\cap\operatorname{\mathcal{B}}=\emptyset$.
+Let $\operatorname{\mathcal{B}}$ be another connected component of $\Psi(A)$.
+Then, there exists an open set $V$ containing $\operatorname{\mathcal{A}}$ such that $V\cap\operatorname{\mathcal{B}}=\emptyset$.
 Recall from Lemma 3 that the two sets $\Psi^{-1}(\operatorname{\mathcal{A}})$ and $\Psi^{-1}(\operatorname{\mathcal{B}})$ are disjoint.
@@
 It follows that $a\in\mathcal{O}$.
-By the assumption, $\Psi$ is upper semi-continuous at $a$.
+By assumption, $\Psi$ is upper semi-continuous at $a$.
 With the open set $V$ defined above, there is an open set $U$ in $\operatorname{dom}\Psi$ such that $\Psi(a^{\prime})\subseteq V$ for all $a^{\prime}\in U$.
-Hence, the openness of $\Psi^{-1}(\operatorname{\mathcal{A}})$ is proved.
-From above, $\Psi^{-1}(\operatorname{\mathcal{A}})$ is both open and closed in the connected set $\operatorname{dom}\Psi$.
+Hence, the openness of $\Psi^{-1}(\operatorname{\mathcal{A}})$ in $\operatorname{dom}\Psi$ is proved.
+From the above, $\Psi^{-1}(\operatorname{\mathcal{A}})$ is both open and closed in the connected set $\operatorname{dom}\Psi$.
 It follows that $\Psi^{-1}(\operatorname{\mathcal{A}})=\operatorname{dom}\Psi$, while $\operatorname{dom}\Psi$ contains $\Psi^{-1}(\operatorname{\mathcal{B}})\neq\emptyset$.
@@
 ∎
-To illustrate above theorem, we consider an example coming from linear fractional vector optimization.
+To illustrate Theorem 4.1, we consider an example coming from linear fractional vector optimization.
 Example 2
@@
 The solution mapping $\Theta_{1}$ of the problem (see Section 5.1) can be written as follows:
-$\Theta_{1}(\xi_{1})=\left\{\begin{array}[]{cl}\{(1,0)\}&\ \ \hbox{ if }\ 0\leq\xi_{1}<\frac{1}{2},\\ [1,+\infty)\times\{0\}&\ \ \hbox{ if }\ \ \ \ \xi_{1}=\frac{1}{2},\\ \left\{\Big{(}\frac{2-3\xi_{1}}{2\xi_{1}-1},2\Big{)}\right\}&\ \ \hbox{ if }\ \frac{1}{2}<\xi_{1}<\frac{2}{3},\\ \{0\}\times[2,+\infty)&\ \ \hbox{ if }\ \ \ \ \xi_{1}=\frac{2}{3},\\ \emptyset&\ \ \hbox{ if }\ \frac{2}{3}<\xi_{1}\leq 1.\\ \end{array}\right.$
+[…]
 The domain of $\Theta_{1}$ is the interval $[0,\frac{2}{3}]$ which is connected.
 Moreover, the set $\mathcal{O}$ is $[0,\frac{1}{2})\cup(\frac{1}{2},\frac{2}{3})$ and $\Theta_{1}$ is single-valued and continuous on $\mathcal{O}$.
-Hence, $\Theta_{1}$ satisfies the condition in above theorem.
+Hence, $\Theta_{1}$ satisfies the condition in Theorem 4.1.
 The image has two connected components, both of them are unbounded.
-The below theorem proposes another criterion to replace connectedness of $\operatorname{dom}\Psi$ in Theorem 4.1.
+The below theorem proposes another criterion to replace the connectedness of $\operatorname{dom}\Psi$ in Theorem 4.1.
 Theorem 4.2
@@
 Note that the openness of $\Psi^{-1}(\operatorname{\mathcal{A}})$ in $\operatorname{dom}\Phi$ is proved under the assumptions that $\mathcal{O}$ is open in $A$ and $A$ is connected.∎
-The assumption related to openness of $\mathcal{O}$ and local boundedness or upper semi-continuity of $\Psi$ on $\mathcal{O}$ in Theorem 4.2 seems heavily in general.
-However, it is satisfied for several cases such as solution mappings in strictly quasi-convex vector optimization (see Lemma 6), linear fractional vector optimization, or monotone affine/general vector inequality yen2011monotone ; hieu2021disconnectedness .
+The assumption related to the openness of $\mathcal{O}$ and local boundedness or upper semi-continuity of $\Psi$ on $\mathcal{O}$ in Theorem 4.2 seems heavy in general.
+However, it is satisfied for several cases such as solution mappings in strictly quasi-convex vector optimization (see Lemma 6), linear fractional vector optimization, or monotone affine/general vector inequalities yen2011monotone ; hieu2021disconnectedness .
 We consider the following example that comes from monotone affine vector inequalities in which the assumption in Theorem 4.1 related to connectedness of the domain is not satisfied but the assumptions in Theorem 4.2 are fulfilled.
@@
 The solution mapping $\Theta_{1}$ of the problem is as follows:
-$\Theta_{1}(\xi_{1})=\left\{\begin{array}[]{cl}\left\{\Big{(}\frac{1}{1-2\xi_{1}},\frac{-1}{1-2\xi_{1}}\Big{)}\right\}&\ \ \hbox{ if }\ 0\leq\xi_{1}<\frac{1}{2}\ \hbox{ or }\ \frac{1}{2}<\xi_{1}\leq 1,\\ \emptyset&\ \ \hbox{ if }\qquad\qquad\ \,\xi_{1}=\frac{1}{2}.\\ \end{array}\right.$
+[…]
 Clearly, $\operatorname{dom}\Theta_{1}=[0,\frac{1}{2})\cup(\frac{1}{2},1]=\mathcal{O}$ is not connected but open in $[0,1]$.
 One observers that $\Theta_{1}$ is single-valued and continuous on $\mathcal{O}$.
-Its image has two connected components, both of them are unbounded.
+Its image has two connected components, both are unbounded.
 Proposition 1
@@
 Example 4
-We consider the mapping given in Example 3.
-The mapping satisfies the condition in Proposition 1.
+We consider the mapping given in Example 3, which satisfies the condition in Proposition 1.
 Clearly, the number of connected components of $\operatorname{dom}\Theta_{1}$ equals the one of the image $\Theta_{1}([0,1])$.
@@
 Let us recall solution concepts and fundamental results in vector optimization.
-Let $K$ be a nonempty, closed and convex subset of $\operatorname{\mathbb{R}}^{n}$ and $f=(f_{1},\dots,f_{m})$ be a continuous mapping from $K$ to $\operatorname{\mathbb{R}}^{m}$.
+Let $K$ be a nonempty and closed subset of $\operatorname{\mathbb{R}}^{n}$ and $f=(f_{1},\dots,f_{m})$ be a continuous mapping from $K$ to $\operatorname{\mathbb{R}}^{m}$.
 The vector minimization problem with the feasible set $K$, vector objective function $f$ and nonnegative orthant +$\operatorname{\mathbb{R}}^{m}_{+}$ as ordering cone is written formally as follows:
@@
 Indeed, $x\in\operatorname{Sol^{w}}(K,f)$ is equivalent to the condition that for every $y\in K$ it is $f(y)-f(x)\not\in-\operatorname{int}\operatorname{\mathbb{R}}^{m}_{+}$, which is the same as $f_{k}(y)\geq f_{k}(x)$ for some index $k$.
-If $f$ is a scalar function, so $f:K\to\operatorname{\mathbb{R}}$, than we point out the fact, that in this case the problem $\operatorname{VOP}(K,f)$ is the well known scalar minimization problem, by writing it as $\operatorname{OP}(K,f)$.
+If $f$ is a scalar function, so $f:K\to\operatorname{\mathbb{R}}$, then we point out the fact, that in this case the problem $\operatorname{VOP}(K,f)$ is the well known scalar minimization problem, by writing it as $\operatorname{OP}(K,f)$.
 In this case the (weak) Pareto solutions are the minimizers of $f$ over $K$.
@@
 $\Delta\coloneqq\{(\xi_{1},\dots,\xi_{m})\in\operatorname{\mathbb{R}}_{+}^{m}:\xi_{1}+\dots+\xi_{m}=1\}.$
-Then its relative interior $\operatorname{ri}\Delta$ is the set of all $\xi$ in $\Delta$ whose coordinates are positive.
+Then, its relative interior $\operatorname{ri}\Delta$ is the set of all $\xi$ in $\Delta$ whose coordinates are positive.
 Clearly, $\Delta$ is nonempty, convex and compact.
@@
 where $\operatorname{Sol}(K,f_{\xi})$ is the solution set of the scalar problem $\operatorname{OP}(K,f_{\xi})$.
-The mapping $\Theta_{1}$ is an effective tool to investigate the weak Pareto solution set and Pareto solution set of a vector optimization problem because of the following results which can be found in, for example, (ehrgott2005multicriteria, , Proposition 3.9) or (ansari2018vector, , Corollaries 4.1–4.3):
+The mapping $\Theta_{1}$ is an effective tool to investigate the weak Pareto solution set and Pareto solution set of a vector optimization problem because of the following results which can be found in, for example, (ehrgott2005multicriteria, , Proposition 3.9) or (ansari2018vector, , Corollaries 4.1, 4.2 & 4.5):
 Lemma 4
 The following inclusions hold for the vector optimization problem $\operatorname{VOP}(K,f)$:
-[…]
+$\Theta_{1}(\Delta)\subseteq\operatorname{Sol^{w}}(K,f)\ \text{and}\ \ \Theta_{1}(\operatorname{ri}\Delta)\subseteq\operatorname{Sol}(K,f).$ (5.2)
+Furthermore, if $K$ is convex and $f$ is convex, then $\Theta_{1}(\Delta)=\operatorname{Sol^{w}}(K,f)$.
 Remark 3
@@
 For this, we consider any $x\in\operatorname{Sol^{w}}(K,f)$.
-If there exists an index $k\in\{1,\dots,m\}$ with $f_{k}(x)=0$ then $x\in\operatorname{Sol}(K,f^{\xi})$, where $\xi=e^{k}$ is the $k$-th standard unit vector of $\operatorname{\mathbb{R}}^{m}$.
+If there exists an index $k\in\{1,\dots,m\}$ with $f_{k}(x)=0$, then $x\in\operatorname{Sol}(K,f^{\xi})$, where $\xi=e^{k}$ is the $k$-th standard unit vector of $\operatorname{\mathbb{R}}^{m}$.
 Indeed, for any $y\in K$ one has $f^{\xi}(y)=\max\{0,f_{k}(y)\}\geq 0$.
-It is clear that $f^{\xi}(x)=0$, this implies $x\in\operatorname{Sol}(K,f^{\xi})$.
+It is clear that $f^{\xi}(x)=0$, which implies $x\in\operatorname{Sol}(K,f^{\xi})$.
 From now on we assume $f(x)\in\operatorname{int}\operatorname{\mathbb{R}}^{m}_{+}$ and define $\xi\in\operatorname{\mathbb{R}}^{m}$ by
 $\xi_{i}:=\frac{1/f_{i}(x)}{1/f_{1}(x)+\dots+1/f_{m}(x)},\quad i=1,\dots,m.$ (5.6)
-Then it is $\xi\in\operatorname{ri}\Delta$ and we get $f^{\xi}(x)=\frac{1}{1/f_{1}(x)+\dots+1/f_{m}(x)}$.
-One has $\xi_{i}f_{i}(x)=\frac{1}{1/f_{1}(x)+\dots+1/f_{m}(x)}$, for all $i\in\{1,\dots,m\}$.
-Because of $x\in\operatorname{Sol^{w}}(K,f)$, for any $y\in K$, there exists an index $k\in\{1,\dots,m\}$ with $f_{k}(y)\geq f_{k}(x)$ implicating
+Then, it is $\xi\in\operatorname{ri}\Delta$ and we get
+$f^{\xi}(x)=\frac{1}{1/f_{1}(x)+\dots+1/f_{m}(x)}.$
+One has
+$\xi_{i}f_{i}(x)=\frac{1}{1/f_{1}(x)+\dots+1/f_{m}(x)},$
+for all $i\in\{1,\dots,m\}$. Because of $x\in\operatorname{Sol^{w}}(K,f)$, for any $y\in K$, there exists an index $k\in\{1,\dots,m\}$ with $f_{k}(y)\geq f_{k}(x)$ implicating
 $f^{\xi}(y)\geq\xi_{k}f_{k}(y)$ $=\frac{f_{k}(y)/f_{k}(x)}{1/f_{1}(x)+\dots+1/f_{m}(x)}$
@@
 which shows $x\in\operatorname{Sol}(K,f^{\xi})$.
-We now prove $\Theta_{2}(\Delta)\subseteq\operatorname{Sol^{w}}(K,f)$.
+We now prove that $\Theta_{2}(\Delta)\subseteq\operatorname{Sol^{w}}(K,f)$.
 Remark 4 points out that $\Theta_{2}(\operatorname{ri}\Delta)\subseteq\operatorname{Sol^{w}}(K,f)$.
-Thus, we need only to show $\Theta_{2}(\partial_{r}\Delta)\subseteq\operatorname{Sol^{w}}(K,f)$.
+Thus, we only need to show $\Theta_{2}(\partial_{r}\Delta)\subseteq\operatorname{Sol^{w}}(K,f)$,
+where $\partial_{r}\Delta$ denotes the relative boundary $\partial_{r}\Delta=\overline{\Delta}\setminus\operatorname{ri}\Delta$.
 Consider $x\in\Theta_{2}(\partial_{r}\Delta)$.
 Again, there exists $\xi\in\partial_{r}\Delta$ such that $x\in\operatorname{Sol}(K,f^{\xi})$.
-If there exists an index $k\in\{1,\dots,m\}$ with $f_{k}(x)=0$ then because of the lower boundedness we immediately get $x\in\operatorname{Sol^{w}}(K,f)$, so that from now on we can assume $f(x)\in\operatorname{int}\operatorname{\mathbb{R}}^{m}_{+}$.
+If there exists an index $k\in\{1,\dots,m\}$ with $f_{k}(x)=0$, then because of the lower boundedness we immediately get $x\in\operatorname{Sol^{w}}(K,f)$, so that from now on we can assume $f(x)\in\operatorname{int}\operatorname{\mathbb{R}}^{m}_{+}$.
 Define
@@
 However, we want to point out that the bound zero is of utmost importance and cannot be generalized without further assumptions.
-In the following, we can see a counter-example.
+In the following, we can see a counter-example in which $\operatorname{Sol^{w}}(K,f)\neq\Theta_{2}(\Delta)$.
 Example 5
@@
 […]
-Figure 3: The relationship $\Theta_{2}(\operatorname{ri}\Delta)\varsubsetneq\operatorname{Sol^{w}}(K,f)\varsubsetneq\Theta_{2}(\operatorname{ri}\Delta)$.
+Figure 3: The relationship $\Theta_{2}(\operatorname{ri}\Delta)\varsubsetneq\operatorname{Sol^{w}}(K,f)\varsubsetneq\Theta_{2}(\Delta)$.
 Remark 6
@@
 Then, the weak Pareto solution set is unbounded.
-Moreover, the Pareto solution set is unbounded provided that $\Theta_{1}(\operatorname{ri}\Delta)$ is nonempty.
+Moreover, the Pareto solution set is unbounded provided that $(\operatorname{ri}\Delta)\cap\operatorname{dom}\Theta_{1}$ is dense in $\operatorname{dom}\Theta_{1}$.
 Proof
-According to Remark 3, the solution mapping $\Theta_{1}$ has closed graph. The conclusions are obtained straightforwardly by applying the conclusions in Theorem 3.1 and inclusions in (5.2). ∎
+According to Remark 3, the solution mapping $\Theta_{1}$ has closed graph. The conclusions are obtained straightforwardly by applying the conclusions from Theorem 3.1 and the inclusions in (5.2). ∎
+Remark 7
+Let $\operatorname{\mathcal{A}}$ be a connected component of $\Theta_{1}(\Delta)$, where $\Theta_{1}$ is given in (5.1).
+If $\Theta_{1}^{-1}(\operatorname{\mathcal{A}})$ is non-closed, then $\operatorname{\mathcal{A}}$ is unbounded.
+Indeed, according to Remark 3, $\Theta_{1}$ has closed graph.
+By applying straightforwardly Remark 1, we obtain the conclusion.
 Example 6
-Consider the bicriteria optimization problem $\operatorname{VOP}(K,f)$ in $\operatorname{\mathbb{R}}^{2}$, with the feasible $K=\operatorname{\mathbb{R}}^{2}$ and two objective functions given by
+Consider the bicriteria optimization problem $\operatorname{VOP}(K,f)$ in $\operatorname{\mathbb{R}}^{2}$, with the feasible set $K=\operatorname{\mathbb{R}}^{2}$ and two objective functions given by
 $f_{1}(x)=x_{1}^{4}-2x_{2},\quad f_{2}(x)=-2x_{1}^{2}+x_{2}^{2}.$
 One can see that the functions are not bounded from below on $K$.
-Thus, two initial points $(0,1),(1,0)$ in $\Delta$ do not belong to the domain of the solution mapping $\Theta_{1}$ of $\operatorname{VOP}(K,f)$.
+Thus, the two end points $(0,1),(1,0)$ in $\Delta$ do not belong to the domain of the solution mapping $\Theta_{1}$ of $\operatorname{VOP}(K,f)$.
 Let $\xi\in\operatorname{ri}\Delta$, equivalently $\xi_{1}\in(0,1)$.
@@
 Since the first two terms in $f_{\xi}$ are nonnegative, $f_{\xi}$ reaches its minimum if and only if these terms are zero.
-Hence, the solution mapping of $\operatorname{VOP}(K,f)$ is as given in (3.1) in Example 1 and its domain is
+Hence, the solution mapping $\Theta_{1}$ of $\operatorname{VOP}(K,f)$ is as given in (3.1) in Example 1 and its domain is
 $\operatorname{dom}\Theta_{1}=\Delta\setminus\{(0,1),(1,0)\}=\operatorname{ri}\Delta.$
@@
 Remark 8
-Let $\operatorname{\mathcal{A}}$ be a connected component of $\Theta_{1}(\Delta)$, where $\Theta_{1}$ is given in (5.1).
-If $\Theta_{1}^{-1}(\operatorname{\mathcal{A}})$ is non-closed, then $\operatorname{\mathcal{A}}$ is unbounded.
-Indeed, according to Remark 3, $\Theta_{1}$ has closed graph.
-By applying straightforwardly Remark 1, we obtain the conclusion.
+By repeating the argument in the proof of Theorem 5.2, we have the same conclusion for the mapping $\Theta_{2}$ given in (5.4).
+If the objective functions are bounded from below by 00 and the domain of the solution mapping $\Theta_{2}$ given in (5.4) is non-closed, then the weak Pareto solution set is unbounded.
 6 Application to quasi-convex vector optimization
@@
 Hence, the solution set of the optimization problem $\operatorname{OP}(K,g)$ is convex.
-In addition, when the inequality in (6.1) is replaced by strict one, we say that $g$ is strictly quasi-convex on $K$.
-Recall that if $g$ is strictly quasi-convex then the solution set of $\operatorname{OP}(K,g)$ has at most one element.
+In addition, when the inequality in (6.1) is replaced by the strict one, we say that $g$ is strictly quasi-convex on $K$.
+Recall that if $g$ is strictly quasi-convex, then the solution set of $\operatorname{OP}(K,g)$ has at most one element.
 6.1 Stability of solutions in parametric optimization under quasi-convexity
@@
 $\Theta(t)\coloneqq\operatorname{Sol}(K,g(\cdot,t)).$ (6.2)
-The following lemma about stability of solutions in quasi-convex optimization plays an important role in the proof of the main result, see Theorem 6.1; Its proof can be found in (bank1983non, , Theorem 4.3.3):
+The following lemma about stability of solutions in quasi-convex optimization plays an important role in this section; its proof can be found in (bank1983non, , Theorem 4.3.3):
 Lemma 5
 Assume that, for any parameter $t\in T$, $g(\cdot,t)$ is quasi-convex on $K$.
-Then, the solution mapping $\Theta$ is upper semicontinuous at all points $t\in T$ such that $\Theta(t)$ is nonempty and bounded.
-Considering the above parametric problem under strict quasi-convexity condition, we have the following lemma which will be used in the proof of Corollary 1.
+Then, the solution mapping $\Theta$ defined in (6.2) is upper semi-continuous at all points $t\in T$ for which $\Theta(t)$ is nonempty and bounded.
+Considering the above parametric problem under the condition of strict quasi-convexity, we have the following lemma which will be used in the proof of Corollary 1.
 Lemma 6
@@
 Proof
-The single-valuedness of $\Theta$ is implied from the strictly quasi-convexity of $g(\cdot,t)$ on $K$.
+The single-valuedness of $\Theta$ is implied by the strict quasi-convexity of $g(\cdot,t)$ on $K$.
 Hence, for arbitrary $t\in\operatorname{dom}\Theta$, $\Theta(t)$ is the unique isolated local solution of $\operatorname{OP}(K,g(\cdot,t))$.
@@
 ∎
+The above lemmas will be used in the proofs of Theorem 6.1, Theorem 6.2, and Theorem 6.3 where the parameter space $T$ is $\Delta$ and the solution mapping $\Theta$ is $\Theta_{1}$ or $\Theta_{2}$.
 6.2 Problems governed by quasi-convex functions
-Under quasi-convexity, to ensure disconnectedness of all components of the weak Pareto solution set, the authors in (hoa2007unbounded, , Theorem 3.2) used a condition which says that the number of components of the set is finite.
+To ensure unboundedness of all components of the weak Pareto solution set of a quasi-convex vector optimization problem, the authors in (hoa2007unbounded, , Theorem 3.2) used the condition that the number of components of the solution set is finite.
 In the following theorem, we propose two other conditions concerning the solution mapping $\Theta_{2}$ given in (5.4) to replace it.
 Let
-$\mathcal{O}=\{\xi\in\operatorname{dom}\Theta_{2}:\Theta_{2}(\xi)\text{ is bounded}\}.$ (6.3)
+$\mathcal{O}_{2}=\{\xi\in\operatorname{dom}\Theta_{2}:\Theta_{2}(\xi)\text{ is bounded}\}.$ (6.3)
 Theorem 6.1
@@
 $\operatorname{dom}\Theta_{2}$ is connected;
-$\mathcal{O}$ is open in $\Delta$.
-If $\operatorname{Sol^{w}}(K,f)$ is disconnected, then all the components are unbounded. Consequently, if the set is bounded then it is connected.
+$\mathcal{O}_{2}$ is open in $\Delta$.
+If $\operatorname{Sol^{w}}(K,f)$ is disconnected, then all its components are unbounded. Consequently, if the set is bounded then it is connected.
 Proof
 Since the objective functions are bounded from below by 00 on $K$, thanks to Theorem 5.1, $\operatorname{Sol^{w}}(K,f)=\Theta_{2}(\Delta)$.
-Moreover, $f_{1},\dots,f_{m}$ are quasi-convex, from (ansari2013generalized, , Theorem 1.31), $f^{\xi}$ given in (5.3) is also quasi-convex on $K$, for $\xi\in\Delta$, and then $\Theta_{2}$ has convex values.
-According to Lemma 5, $\Theta_{2}$ is upper semi-continuous on $\mathcal{O}$.
+Moreover, $f_{1},\dots,f_{m}$ are quasi-convex, from (ansari2013generalized, , Theorem 1.31), $f^{\xi}$ given in (5.3) is also quasi-convex on $K$, for $\xi\in\Delta$, and, then $\Theta_{2}$ has convex values.
+According to Lemma 5, $\Theta_{2}$ is upper semi-continuous on $\mathcal{O}_{2}$.
 Therefore, the desired conclusion is obtained by applying directly Theorems 4.1 and 4.2 for $\Theta_{2}$ with the note that $\Theta_{2}$ has closed graph and connected values.
@@
 Suppose that $\Theta_{2}$ is single-valued on its domain.
-If $\operatorname{Sol}(K,f)$ is disconnected then all its components are unbounded.
+If $\operatorname{Sol}(K,f)$ is disconnected, then all its components are unbounded.
 Proof
-From the assumption, the domain of $\Theta_{2}$ is $\mathcal{O}$, where $\mathcal{O}$ is defined by (6.3).
-Hence, $\mathcal{O}$ is the set of $\xi$ in $\Delta$ such that $\Theta_{2}(\xi)$ has only one element.
+From the assumption, the domain of $\Theta_{2}$ is $\mathcal{O}_{2}$, where $\mathcal{O}_{2}$ is defined by (6.3).
+Hence, $\mathcal{O}_{2}$ is the set of $\xi$ in $\Delta$ such that $\Theta_{2}(\xi)$ has only one element.
 Thanks to (warburton1983quasiconcave, , Theorem 3.5), we claim that $\operatorname{Sol}(K,f)$ coincides with $\operatorname{Sol^{w}}(K,f)$.
@@
 Both $f_{1},f_{2}$ are strictly bounded from below by 00 on $K$. This is a quasi-convex problem and its solution mapping $\Theta_{2}$ is given by
-$\Theta_{2}(\xi_{1})=\left\{\begin{array}[]{cl}\emptyset&\ \ \hbox{ if }\ 0\leq\xi_{1}<\frac{1}{2}\hbox{ or }\ \frac{5}{6}<\xi_{1}\leq 1,\\ \left\{\Big{(}\frac{1+2\xi_{1}}{1-2\xi_{1}},0\Big{)}\right\}&\ \ \hbox{ if }\qquad\quad\ \frac{1}{6}\leq\xi_{1}<\frac{1}{2},\\ \left\{\Big{(}\frac{3-2\xi_{1}}{2\xi_{1}-1},4\Big{)}\right\}&\ \ \hbox{ if }\qquad\quad\ \frac{1}{2}<\xi_{1}\leq\frac{5}{6}.\end{array}\right.$
+[…]
 Clearly, the domain of $\Theta_{2}$ is the union of two open intervals $\left(\frac{1}{6},\frac{1}{2}\right)\cup\left(\frac{1}{2},\frac{5}{6}\right)$; it is disconnected.
 The mapping is single-valued on its domain.
-According to Corollary 1, the Pareto solution set and weak Pareto solution set are coincident.
-They are the union of two disconnected parallel lines:
+According to Corollary 1, the Pareto solution set and weak Pareto solution set coincide.
+They are the union of two disconnected parallel rays:
 $\{(x_{1},x_{2})\in\operatorname{\mathbb{R}}^{2}:x_{1}\geq 2,x_{2}=0\}\cup\{(x_{1},x_{2})\in\operatorname{\mathbb{R}}^{2}:x_{1}\geq 2,x_{2}=4\}.$
-Remark 8
-If all $f_{i}$ are convex, we have the same conclusions in Theorem 6.1 for the solution map $\Theta_{1}$ defined by (5.1).
-However, until now, we have not known any explicit convex vector optimization problems whose (weak) Pareto solution set is disconnected.
-This is related to Question 9.8 in yen2011monotone and Question 5.1 in hieu2021disconnectedness .
-In the case that all objective functions are strictly quasi-convex, $\Theta_{2}$ is single-valued on its domain. Therefore, we have the following theorem:
-Theorem 6.2
+In the case that all objective functions are strictly quasi-convex, $\Theta_{2}$ is single-valued on its domain. Therefore, we have the following result:
+Corollary 2
 Suppose that all $f_{i}$ are strictly quasi-convex and strictly bounded from below by 00 on $K$.
 Then, one has $\operatorname{Sol}(K,f)=\operatorname{Sol^{w}}(K,f)$.
-Moreover, if $\operatorname{Sol}(K,f)$ is disconnected then all its components are unbounded.
+Moreover, if $\operatorname{Sol}(K,f)$ is disconnected, then all its components are unbounded.
 Proof
 For $\xi\in\Delta$, $f^{\xi}$ is also strictly quasi-convex on $K$ (see (warburton1983quasiconcave, , Lemma 4.2)); Hence $\Theta_{2}$ is single-valued on its domain.
-The conclusions are implied straightforwardly from Corollary 1.
+The conclusions are implied straightforwardly by Corollary 1.
 ∎
+In the following theorem, we provide a necessary and sufficient condition for unboundedness of the weak Pareto solution set of $\operatorname{VOP}(K,f)$ under certain conditions:
+Theorem 6.2
+Let $f_{1},\dots,f_{m}$ be quasi-convex and bounded from below by 00 on $K$.
+Assume that $\Theta_{2}$ has bounded values.
+Then, the weak Pareto solution set is unbounded if and only if $\operatorname{dom}\Theta_{2}$ is non-closed in $\Delta$.
+Proof
+According to Theorem 5.1, one has $\operatorname{Sol^{w}}(K,f)=\Theta_{2}(\Delta)$.
+Clearly, if $\operatorname{dom}\Theta_{2}$ is non-closed in $\Delta$, thanks to Remark 8, the weak Pareto solution set is unbounded.
+We now prove the inverse conclusion.
+Assume that $\operatorname{Sol^{w}}(K,f)$ is unbounded, but on the contrary, $\operatorname{dom}\Theta_{2}$ is closed in $\Delta$.
+Let $\{x^{k}\}$ be an unbounded sequence in $\operatorname{Sol^{w}}(K,f)$.
+Because of $\operatorname{Sol^{w}}(K,f)=\Theta_{2}(\Delta)$, there exist $\{\xi^{k}\}\subset\operatorname{dom}\Theta_{2}\subset\Delta$ such that $x^{k}\in\operatorname{Sol}(K,f_{\xi^{k}})$, for each $k$.
+By boundedness and closedness of $\operatorname{dom}\Theta_{2}$, we assume that $\xi^{k}\to\bar{\xi}\in\operatorname{dom}\Theta_{2}$, i.e., $\Theta_{2}(\bar{\xi})\neq\emptyset$.
+Since $\Theta_{2}$ has bounded values, $\Theta_{2}(\bar{\xi})$ is bounded.
+Let $O$ be an open and bounded ball containing $\Theta_{2}(\bar{\xi})$.
+Thanks to Lemma 5, $\Theta_{2}$ is upper semi-continuous at $\bar{\xi}$, hence there exists an open neighbourhood $\Delta_{O}$ of $\bar{\xi}$ such that $\Theta_{2}(\xi)\subset O$ for all $\xi\in\Delta_{O}$.
+This implies that $x^{k}\in O$ for all $k$ large enough.
+This is impossible since $\{x^{k}\}$ is unbounded.
+∎
+6.3 Problems governed by convex functions
+For problems given by convex data, we also have a necessary and sufficient condition for unboundedness of the weak Pareto solution set of $\operatorname{VOP}(K,f)$ as follows:
+Theorem 6.3
+Let $K$ be convex and $f_{1},\dots,f_{m}$ be convex over $K$.
+Assume that $\Theta_{1}$ has bounded values.
+Then, the weak Pareto solution set is unbounded if and only if $\operatorname{dom}\Phi_{1}$ is non-closed in $\Delta$.
+Proof
+We consider the scalarization $f_{\xi}=\xi_{1}f_{1}+\dots+\xi_{m}f_{m}$, for $\xi\in\Delta$.
+Clearly, $f_{\xi}$ is also convex over $K$.
+Moreover, according to Lemma 4, one has $\operatorname{Sol^{w}}(K,f)=\Phi_{1}(\Delta)$.
+The conclusion is proved by repeating the argument in the proof of Theorem 6.2.
+∎
+Remark 9
+If $K$ is convex and all $f_{i}$ are convex over $K$, we have the same conclusions in Theorem 6.1 for the solution map $\Theta_{1}$ defined by (5.1).
+However, until now, we still do not know any explicit convex vector optimization problems whose (weak) Pareto solution set is disconnected.
+This is related to Question 9.8 in yen2011monotone and Question 5.1 in hieu2021disconnectedness .
+When $\operatorname{VOP}(K,f)$ is given by convex polynomial data, i.e. $f_{1},\dots,f_{m}$ are polynomials and $K$ is given by convex polynomials as follows:
+$K=\{x\in\operatorname{\mathbb{R}}^{n}:g_{1}(x)\leq 0,\dots,g_{s}(x)\leq 0\},$
+under a mild condition, in the following theorem, we claim that $\operatorname{Sol^{w}}(K,f)$ is connected.
+Theorem 6.4
+Let $f_{1},\dots,f_{m}$ and $g_{1},\dots,g_{s}$ be polynomial and convex. Assume that $\Theta_{1}$ has bounded values. Then, the weak Pareto solution set is connected.
+Proof
+Recall that $f_{\xi}$ is also convex over $K$ and $\operatorname{Sol^{w}}(K,f)=\Phi_{1}(\Delta)$ (see Lemma 4).
+Assume that $\operatorname{Sol^{w}}(K,f)$ is nonempty, this implies that $\operatorname{dom}\Phi_{1}$ is so.
+We prove that $\operatorname{dom}\Phi_{1}$ is convex.
+Indeed, let $\xi^{\prime},\xi^{\prime\prime}$ be two points in $\operatorname{dom}\Phi_{1}$, i.e., $f_{\xi^{\prime}}$ and $f_{\xi^{\prime\prime}}$ are bounded from below over $K$.
+This implies that $f_{\xi}$ is also bounded from below over $K$ for all $\xi$ in the line segment connecting $\xi^{\prime}$ and $\xi^{\prime\prime}$, denoted by $[\xi^{\prime},\xi^{\prime\prime}]$.
+According to the Frank–Wolfe type theorem in convex polynomial optimization belousov2002frank , $\operatorname{Sol}(K,f_{\xi})$ is nonempty for all $\xi\in[\xi^{\prime},\xi^{\prime\prime}]$.
+Hence, $\operatorname{dom}\Phi_{1}$ is convex.
+Therefore, connectedness of $\operatorname{Sol^{w}}(K,f)$ is straightforward from Lemma 5 and Theorem 2.1. ∎
+To end this section, we illustrate Theorems 6.3 and 6.4 by considering the following example:
+Example 8
+Consider the convex polynomial bicriteria optimization problem $\operatorname{VOP}(K,f)$ in $\operatorname{\mathbb{R}}^{2}$, with the convex feasible set $K=\operatorname{\mathbb{R}}_{+}^{2}$ and two convex objective functions given by
+$f_{1}(x)=x_{1}^{2}-x_{2},\quad f_{2}(x)=-x_{1}+x_{2}^{2}.$
+Clearly, these functions are not bounded from below on $K$.
+Thus, the two end points $(0,1),(1,0)$ in $\Delta$ do not belong to the domain of the solution mapping $\Theta_{1}$ of $\operatorname{VOP}(K,f)$.
+Let $\xi\in\operatorname{ri}\Delta$, equivalently $\xi_{1}\in(0,1)$.
+One has
+$f_{\xi}=\xi_{1}f_{1}+(1-\xi_{1})f_{2}=\xi_{1}x_{1}^{2}+(1-\xi_{1})x_{2}^{2}-(1-\xi_{1})x_{1}-\xi_{1}x_{2},$
+which is quadratic and convex over $K$. An easy computation shows that $\operatorname{Sol}(K,f_{\xi})$ has the unique solution
+$\operatorname{Sol}(K,f_{\xi})=\Big{\{}\Big{(}\frac{1-\xi_{1}}{2\xi_{1}},\frac{\xi_{1}}{2(1-\xi_{1})}\Big{)}\Big{\}},\,\xi_{1}\in(0,1).$
+One sees that the weak Pareto solution set is connected, moreover, the set is unbounded and $\Theta_{1}$ has non-closed domain.
 Perspectives and future work
-The results related to set-valued mappings having connected values in this paper can be useful for investigating (un)bounedness and (un)connectedness of solution sets in set optimization, vector equilibrium, or vector variational inequality under some generalized convexity/monotone assumptions.
-We plan to exploit special structures such as polyhedral or semi-algebraic to discover new properties concerning contractibility or connectedness of solution sets of these problems as in huong2017connectedness ; huong2016polynomial ; hieu2019polynomialVVIS ; hieu2020application .
+The results related to set-valued mappings having connected values in this paper can be useful for investigating (un)bounedness and (dis)connectedness of solution sets in set optimization, vector equilibrium, or vector variational inequality under some generalized convexity/monotone assumptions.
+We plan to exploit special structures in geometry such as polyhedral or semi-algebraic to discover new properties concerning contractibility or connectedness of solution sets of these problems as in huong2017connectedness ; huong2016polynomial ; hieu2019polynomialVVIS ; hieu2020application .
+Acknowledgements.
+The authors are indebted to the anonymous referee for the insightful comments and valuable suggestions. The ERCIM Alain Bensoussan Fellowship Programme supported the research of Vu Trung Hieu.